### v1

# Imaging heat transport in suspended diamond nanostructures with integrated spin defect thermometers

Abstract
Among all materials, mono-crystalline diamond has one of the highest measured thermal conductivities, with values above 2000 W/m/K at room temperature. This stems from momentum-conserving ‘normal’ phonon-phonon scattering processes dominating over momentum-dissipating ‘umklapp’ processes, a feature that also suggests diamond as an ideal platform to experimentally investigate phonon heat transport phenomena that violate Fourier’s law. Here, we introduce dilute nitrogen-vacancy colour centers as in-situ, highly precise spin defect thermometers to image temperature inhomogeneities in single-crystal diamond microstructures heated from ambient conditions. We analyse cantilevers with cross-sections in the range from about 0.2 to 2.6 m2, observing a relation between cross-section and heat flux that departs from Fourier’s law predictions. We rationalize such behavior relying on first-principles simulations based on the linearized phonon Boltzmann transport equation, also discussing how fabrication-induced impurities influence conduction. Our temperature-imaging method can be applied to diamond devices of arbitrary geometry, paving the way for the exploration of unconventional, non-diffusive heat transport phenomena.
Carbon-based materials such as graphene, graphite or diamond present exceptional heat-conduction properties. Due to large phonon group velocities, they exhibit high thermal conductivity – with diamond reaching record values at room temperature and below [1, 2, 3]. Moreover, the prevalence of normal (momentum-conserving) phonon-phonon scattering processes over Umklapp (momentum non-conserving) processes in these materials leads to spectacular phonon heat transport phenomena, that violate Fourier’s law for macroscopic diffusive heat propagation [4]. In graphene and graphite, ballistic transport in nanoscale devices [5, 6] or hydrodynamic heat transport [7, 8, 9, 10, 11] have been evidenced at or near room temperature. While recent theoretical work predicted that diamond should also exhibit heat transport deviating from Fourier’s diffusion at room temperature [12], little experimental work has been reported in this direction so far. Noteworthy, the influence of phonon scattering at grain boundaries has been explored in polycrystalline diamond [13] and unexpected size effects were recently reported in the thermal conductivity of diamond microparticles [14]. Understanding heat transport in nanostructured diamond is also valuable in view of its applications in optomechanics [15, 16], nonlinear optics [17], quantum nanophotonics [18], as well as in heat management devices [19].
A major challenge in the study of thermal transport at the nanoscale is to find appropriate temperature sensors. Traditional measurements of the heat conductivity rely on resistive sensors combined with advanced sample engineering [5, 9] and cannot provide spatially-resolved temperature measurements. Time-resolved non contact optical thermometry techniques such as transient thermal gratings [20, 7], and time- or frequency domain thermoreflectance [21, 22, 23] have excellent temporal resolution but limited spatial resolution, and are often applicable to simple planar geometries only. Other optical techniques [24, 25], including Raman sideband thermometry [26, 27, 28], offer good spatial resolution but rather poor temperature accuracy, typically not better than a few K.
Fittingly, diamond hosts optically active spin defects, the most studied being the negatively charged nitrogen vacancy (NV) center [29], which can be used as in-situ nanoscale thermometers [30, 31, 32]. Negatively charged NV centers are spin-1 atomic defects formed by a substitutional nitrogen atom neighboring a vacancy in the diamond lattice. Their spin can be initialized and readout by optical pumping and monitoring the photoluminescence intensity. The resonance frequency associated with the transition between and spin states is temperature-dependent [33], as can be tracked with either continuous-wave or pulsed optically detected magnetic resonance (ODMR) [34]. NV centers are thus excellent thermometers, offering nanoscale spatial resolution and sub- sensitivity [35] for steady-state temperature measurement near room temperature. Solid-state spin defects acting as temperature sensors are also found in a broad range of materials, including hBN [36], Si [37], SiC [38, 39, 40] and this field is currently undergoing rapid developments. However, the potential of quantum sensors in the study of non-Fourier thermal transport remains largely unexplored and unexploited.
In this work, we use dilute NV centers in the diamond lattice as precise in-situ temperature sensors to investigate heat transport properties of suspended diamond microstructures. The principle of our experiment is summarized in Fig.1a. We fabricate suspended micro-cantilevers out of single crystal diamond grown by chemical vapour deposition and enriched in (%), containing NV centers up to a concentration of 3 ppm. Cantilevers are patterned using electron-beam lithography and subsequent etching, in two steps: first a step of vertical, oxygen-based reactive ion etching, followed by focused ion beam (FIB) milling at an angle from the diamond surface. This results in triangular cantilever cross-sections, as visible in Fig. 1b. The cantilever acts as a one-dimensional channel for heat transport. A small chromium patch is added at the cantilever tip to serve as a heat source, activated by optical absorption of a 515 nm laser focused onto it, referred to as heating laser. The bulk diamond acts as a heat sink. Complete details on the sample fabrication can be found in the Supplementary Material (SM) [41].
Photoluminescence from NV centers is collected in confocal geometry, using a second 515 nm laser beam to address a large ensemble of NV centers (about ) located in the diffraction-limited spot. This second laser is referred to as readout laser. The temperature-dependent zero-field splitting between and states is measured through continuous wave (cw) ODMR. We use a lock-in amplifier (LIA) combined with frequency modulation of the microwave to achieve high signal to noise ratio [42]. Fig. 1c presents the shift of ZFS value versus temperature, defined as , measured both in bulk diamond and on cantilevers. Note that, at room temperature, the ZFS value is around 2.87 GHz. A position-dependent offset can be observed, due for example to the presence of local strain, particularly in cantilevers. As visible in Fig. 1c, this offset does not modify the temperature-dependent relative ZFS shift. Our data is matching previous reports, and the ZFS shift is fitted by the model from Ref. [43], with only two fitting parameters (dashed line in Fig. 1c).
The calibration is used to convert values of ZFS shift to local temperature under the readout laser spot. Noise analysis gives a typical temperature sensitivity of 0.27 for the temperature readout on the cantilever [41]. Note that the temperature sensitivity depends on the number of NV probed by the readout spot, it can thus degrade when addressing smaller cantilevers, but also vary with the NV density in the starting diamond substrate. The choice of cw-ODMR is a trade-off between sensitivity requirement and simplicity of experimental implementation. Even though more sophisticated NV-based thermometry schemes allow to reach better sensitivity, relying e.g. on pulse sequences [30, 31] or on multi-frequency drive [35], in our case uncertainties in the absorbed laser power and exact cantilever cross-sections have larger impacts on the final error bars, as detailed below.
The heating laser is then used to induce a heat flow in a cantilever. Fig. 2a presents the measured local temperature shift as a function of heating laser power , for a readout spot located at a fixed position, in the middle of the cantilever. We observe a linear relation between and heating power. Standard deviations of residuals after a linear fit give a temperature error of K [41]. This value includes systematic errors due e.g to imperfect alignment of the heating laser on the gold patch, explaining why it is slightly higher than the readout error ( mK with our 30 s integration time).
Fixing the heating laser power, we measure the space-dependent temperature profile along the cantilever by sweeping the position of the readout laser. For each readout laser position, two measurements are necessary to extract the local heating-induced temperature shift: is obtained by subtracting the ZFS value without and with heating. The temperature profile for a cantilever that has a narrower section in the middle, see Fig. 2b, is shown in Fig. 2c. A section of constant width shows linear temperature profile. Moreover, the narrow section has a much steeper temperature gradient than the wide section, while the temperature is constant on the free-hanging side, away from the Cr patch. In our experimental setting, the spatial resolution of the temperature measurement is diffraction-limited, the readout laser spot size being 0.44 µm.
We emphasize that the ZFS can also shift due to local strain or stress. This is for example visible in the evolution of ZFS along the cantilever (in the absence of heating), where a strong variation is observed near the cantilever base [41]. However, since our cantilevers are free-standing, we do not expect that heating induces any additional strain except near its base. We also confirmed that the Cr patch is well thermalized with the diamond by independent estimation of its temperature using anti-Stokes emission spectroscopy [41].
We can consider the case of standard diffusive heat transport to develop a picture of the behavior expected from the established Fourier’s law. According to Fourier’s law, in 1D, , with the local heat flux density and the material’s thermal conductivity. Neglecting all other thermal losses (mainly convective or radiative), Fourier’s law predicts a constant heat flux along the cantilever , with the heating laser power incident on the Cr patch, the Cr patch absorption coefficient and the cantilever cross-section. is temperature dependent, but for small heating power, Fourier’s law predicts a linear temperature profile along the cantilever, with a local temperature shift directly proportional to the heating laser power, compatible with the results from Fig. 2. However, having a linear temperature profile is a necessary but not sufficient condition for the validity of Fourier’s law. In fact, Fourier’s law also assumes thermal conductivity to be an intrinsic, size-independent material’s property, and consequently a heat flux that scales linearly with the cantilever cross-section. On the contrary, when reducing the lateral size of a 1D channel, we intuitively expect size effect to become increasingly more relevant, leading to a reduction of . In the following we investigate the heat flux’s magnitude as a function of cantilever width.
In diamond, has been predicted to decreases to approximately half of its bulk value at widths 0.7m [44, 45] – a much stronger reduction than other materials for this channel size[46, 47, 48, 49, 50]. However, the width dependence of thermal conductivity in diamond has not been experimentally demonstrated yet. Our experimental platform is ideally suited for investigating this phenomenon. We investigate cantilevers with different lateral width , in the range 0.8-2.5 µm, and identical length of 10 µm. The lower bound of this range originates from limitations in temperature sensitivity for cantilever widths below the diffraction limited spot size (due to the reduced number of probed NV centers). Conversely, the upper bound is due to difficulty in fabricating suspended cantilevers with dimensions bigger than 3 µm (due to the long etching time required). We measure the temperature profile along each individual cantilever, adjusting such that the total temperature shift is comparable in all cases and remains below 40 K. The results are presented in Fig. 3, with panel c presenting the measured temperature gradient , normalized by , versus cantilever width . In the case of purely diffusive thermal transport, Fourier law predicts that this normalized temperature gradient is inversely proportional to the cantilevers cross section. Specifically, since all our cantilevers have triangular cross section with same angle , Fourier’s law predicts . In contrast, our experiments deviate from Fourier’s prediction, with a temperature gradient displaying a scaling approximatively proportional to for µm. This indicates that the probed range of cantilever width is within the non-Fourier regime, where non-diffusive effects are relevant and thermal conductivity depends on the sample’s geometry (size).
We model thermal transport in our cantilevers using the viscous heat equations (VHE), two coupled mesoscopic heat-transport equations for temperature and phonon drift velocity [12] that generalize Fourier’s laws encompassing not only heat diffusion but also fluid-like behavior [7, 8, 9, 10, 51]. The VHE can be parametrized from first principles and allow to describe how micro-structured geometries influence thermal transport with accuracy comparable to microscopic models based on the full space-dependent solution of the Linearized Boltzmann transport equation (LBTE) [52, 53], at much reduced computational cost [54]. Therefore, we employ the VHE to investigate how the geometry of our cantilevers affects thermal transport, and we assess the relevance of viscous effects by comparing the VHE predictions with those obtained from Fourier’s law.
We simulate the three-dimensional experimental geometry, imposing a temperature gradient K around room temperature on µm long cantilevers. Parameters entering in the VHE or Fourier’s law are computed from first-principles, accounting for third-order anharmonic phonon-phonon scattering [55, 45], phonon-impurity scattering [56], as well as for the effects of rough cantilever’s boundaries [44] (see SM [41] for details). Then, we solve the VHE or Fourier’s law and determine the heat flux from their solution. Finally, we evaluate the effective thermal conductivity as , where is the cantilever cross section, and the oriented surface element.
In Fig. 4 we compare the experimental values for , extracted from the measurements in Fig. 3 (see SM [41] for details), against predictions obtained solving the VHE or Fourier’s law. In the numerical simulations, we consider two different impurity scenarios to compute transport coefficients from first principles. First, we consider the ideal case of pristine diamond cantilevers with only 0.05% 13C isotopes (dashed lines). In this idealized case, VHE and Fourier’s law yield values having practically unimportant differences, however both significantly higher than the experimental values.
To address this discrepancy, we investigate theoretically the effect of possible additional sources of thermal resistance missing from our simulations in pristine samples. In particular, the cantilever’s fabrication procedure based on FIB milling is known to cause contamination of Ga ions in proximity of the milled surfaces [57] (Fig. 4 inset). Ga contamination can be as high as hundreds of ppm within a penetration layer of depth estimated in the present case to µm [41]. The penetration layer is determined by the FIB specifications, it is therefore the same for cantilevers having different cross sections. This implies that the average concentration of Ga impurities in our samples is proportional to the surface-to-volume ratio. Consequently, Ga contamination is expected to have an increasingly more important effect as the cantilevers width decreases. Therefore, we account for the influence of Ga impurities using the model by Ref. [56], considering an average concentration of Ga impurities proportional to the surface-to-volume ratio, and determine one parameter — the concentration of Ga impurities in the penetration layer — from a fit with experimental data [41]. NV center impurities also present in our diamond samples are negelected due to their small concentration (3 ppm) [58].
Solid lines in Fig. 4 are computed values for obtained from this one-parameter model, where the concentration of Ga impurities in the penetration layer obtained from the fit is ppm (a value of the same order of magnitude of those discussed in Ref. [57]). We highlight the good quantitative agreement between predictions and experiments. Contrarily, a model with constant concentration of impurities over the whole cross-section fails to reproduce the experiments, especially in the trend [41]. This suggests that the main hypothesis of our model — having average concentration of Ga impurities proportional to the surface-to-volume ratio of the cantilevers — is realistic and captures the most important microscopic physics. Finally, hydrodynamic effects (quantified by the difference between green and red curves) appear practically unimportant and the deviation from Fourier’s law are mainly caused by being not an intrinsic property but a size-dependent quantity.
In conclusion, we highlight that this platform offers multiple routes for new investigations. First, the question arises of how to maximize heat conductivity in sub-micrometer diamond channels, which can be answered by achieving a finer control on the impurity level. The effect of sidewall roughness on phonon scattering at the boundaries [59, 60] can also be probed by varying the etching method and chemical termination of the diamond surface. Second, hydrodynamic effects in heat transport could be revealed using refined geometries that enhance unambiguous signatures of hydrodynamic features, such as heat vortices [54]. Third, this research can be pushed toward even smaller channels to address open questions in quantized phononic heat conductance [61]. Overall, our results demonstrate the potential of solid state spin defects in ultrahigh thermal conductivity materials to explore heat transport phenomena in non-Fourier regimes.
Acknowledgements.
Acknowledgments. This project has received funding from the Swiss National Science Foundation through grant #198898 and #216406, EPFL Center for Quantum Science and Engineering through a Collaborative Research Fellowship, the European Union’s Horizon 2020 research and innovation programmes under the Marie Skłodowska-Curie grant agreement #945363 and under grant agreement #101112347, Project NerveRepack. M. S. acknowledges support from: (i) the Kelvin2 HPC platform at the NI-HPC Centre (funded by EPSRC and jointly managed by Queen’s University Belfast and Ulster University); (ii) the UK National Supercomputing Service ARCHER2, for which access was obtained via the UKCP consortium and funded by EPSRC [EP/X035891/1]; (iii) Gonville and Caius College. N.M and E.D. acknowledge support from the Swiss National Science Foundation (SNSF), through Grant No. CRSII5_189924 (“Hydronics” project). The diamond cantilevers were fabricated at the EPFL Center of MicroNanoTechnology (CMi) and we gratefully acknowledge the support of EPFL CMi.References
- Berman et al. [1975] R. Berman, P. R. W. Hudson, and M. Martinez, Nitrogen in diamond: evidence from thermal conductivity, Journal of Physics C: Solid State Physics 8, L430 (1975).
- Olson et al. [1993] J. R. Olson, R. O. Pohl, J. W. Vandersande, A. Zoltan, T. R. Anthony, and W. F. Banholzer, Thermal conductivity of diamond between 170 and 1200 k and the isotope effect, Phys. Rev. B 47, 14850 (1993).
- Inyushkin et al. [2018] A. V. Inyushkin, A. N. Taldenkov, V. G. Ralchenko, A. P. Bolshakov, A. V. Koliadin, and A. N. Katrusha, Thermal conductivity of high purity synthetic single crystal diamonds, Phys. Rev. B 97, 144305 (2018).
- Chen [2021] G. Chen, Non-fourier phonon heat conduction at the microscale and nanoscale, Nature Reviews Physics , 1 (2021).
- Bae et al. [2013] M.-H. Bae, Z. Li, Z. Aksamija, P. N. Martin, F. Xiong, Z.-Y. Ong, I. Knezevic, and E. Pop, Ballistic to diffusive crossover of heat flow in graphene ribbons, Nature Communications 4, 1734 (2013).
- Nika and Balandin [2012] D. L. Nika and A. A. Balandin, Two-dimensional phonon transport in graphene, Journal of Physics: Condensed Matter 24, 233203 (2012).
- Huberman et al. [2019] S. Huberman, R. A. Duncan, K. Chen, B. Song, V. Chiloyan, Z. Ding, A. A. Maznev, G. Chen, and K. A. Nelson, Observation of second sound in graphite at temperatures above 100 k, Science 364, 375 (2019).
- Ding et al. [2022] Z. Ding, K. Chen, B. Song, J. Shin, A. A. Maznev, K. A. Nelson, and G. Chen, Observation of second sound in graphite over 200 k, Nature Communications 13, 285 (2022).
- Machida et al. [2020] Y. Machida, N. Matsumoto, T. Isono, and K. Behnia, Phonon hydrodynamics and ultrahigh–room-temperature thermal conductivity in thin graphite, Science 367, 309 (2020).
- Jeong et al. [2021] J. Jeong, X. Li, S. Lee, L. Shi, and Y. Wang, Transient hydrodynamic lattice cooling by picosecond laser irradiation of graphite, Phys. Rev. Lett. 127, 085901 (2021).
- Huang et al. [2023] X. Huang, Y. Guo, Y. Wu, S. Masubuchi, K. Watanabe, T. Taniguchi, Z. Zhang, S. Volz, T. Machida, and M. Nomura, Observation of phonon poiseuille flow in isotopically purified graphite ribbons, Nature Communications 14, 2044 (2023).
- Simoncelli et al. [2020] M. Simoncelli, N. Marzari, and A. Cepellotti, Generalization of fourier’s law into viscous heat equations, Phys. Rev. X 10, 011019 (2020).
- Sood et al. [2018] A. Sood, R. Cheaito, T. Bai, H. Kwon, Y. Wang, C. Li, L. Yates, T. Bougher, S. Graham, M. Asheghi, M. Goorsky, and K. E. Goodson, Direct Visualization of Thermal Conductivity Suppression Due to Enhanced Phonon Scattering Near Individual Grain Boundaries, Nano Letters 18, 3466 (2018).
- Wang and Sun [2024] Y. Wang and B. Sun, Anomalously strong size effect on thermal conductivity of diamond microparticles, Applied Physics Letters 125, 042202 (2024).
- Tao et al. [2014] Y. Tao, J. M. Boss, B. A. Moores, and C. L. Degen, Single-crystal diamond nanomechanical resonators with quality factors exceeding one million, Nature Communications 5, 3638 (2014).
- Mitchell et al. [2016] M. Mitchell, B. Khanaliloo, D. P. Lake, T. Masuda, J. P. Hadden, and P. E. Barclay, Single-crystal diamond low-dissipation cavity optomechanics, Optica 3, 963 (2016).
- Hausmann et al. [2014] B. J. M. Hausmann, I. Bulu, V. Venkataraman, P. Deotare, and M. Lončar, Diamond nonlinear photonics, Nature Photon 8, 369 (2014).
- Moody et al. [2022] G. Moody, V. J. Sorger, D. J. Blumenthal, P. W. Juodawlkis, W. Loh, C. Sorace-Agaskar, A. E. Jones, K. C. Balram, J. C. F. Matthews, A. Laing, M. Davanco, L. Chang, J. E. Bowers, N. Quack, C. Galland, I. Aharonovich, M. A. Wolff, C. Schuck, N. Sinclair, M. Lončar, T. Komljenovic, D. Weld, S. Mookherjea, S. Buckley, M. Radulaski, S. Reitzenstein, B. Pingault, B. Machielse, D. Mukhopadhyay, A. Akimov, A. Zheltikov, G. S. Agarwal, K. Srinivasan, J. Lu, H. X. Tang, W. Jiang, T. P. McKenna, A. H. Safavi-Naeini, S. Steinhauer, A. W. Elshaari, V. Zwiller, P. S. Davids, N. Martinez, M. Gehl, J. Chiaverini, K. K. Mehta, J. Romero, N. B. Lingaraju, A. M. Weiner, D. Peace, R. Cernansky, M. Lobino, E. Diamanti, L. T. Vidarte, and R. M. Camacho, 2022 roadmap on integrated quantum photonics, Journal of Physics: Photonics 4, 012501 (2022).
- Liu et al. [2015] X. Liu, G. Zhang, and Y.-W. Zhang, Surface-engineered nanoscale diamond films enable remarkable enhancement in thermal conductivity and anisotropy, Carbon 94, 760 (2015).
- Minnich et al. [2011] A. J. Minnich, J. A. Johnson, A. J. Schmidt, K. Esfarjani, M. S. Dresselhaus, K. A. Nelson, and G. Chen, Thermal conductivity spectroscopy technique to measure phonon mean free paths, Phys. Rev. Lett. 107, 095901 (2011).
- Regner et al. [2013] K. T. Regner, D. P. Sellan, Z. Su, C. H. Amon, A. J. H. McGaughey, and J. A. Malen, Broadband phonon mean free path contributions to thermal conductivity measured using frequency domain thermoreflectance, Nature Communications 4, 1640 (2013).
- Hu et al. [2015] Y. Hu, L. Zeng, A. J. Minnich, M. S. Dresselhaus, and G. Chen, Spectral mapping of thermal conductivity through nanoscale ballistic transport, Nature Nanotechnology 10, 701 (2015).
- Tian et al. [2018] F. Tian, B. Song, X. Chen, N. K. Ravichandran, Y. Lv, K. Chen, S. Sullivan, J. Kim, Y. Zhou, T.-H. Liu, M. Goni, Z. Ding, J. Sun, G. A. G. Udalamatta Gamage, H. Sun, H. Ziyaee, S. Huyan, L. Deng, J. Zhou, A. J. Schmidt, S. Chen, C.-W. Chu, P. Y. Huang, D. Broido, L. Shi, G. Chen, and Z. Ren, Unusual high thermal conductivity in boron arsenide bulk crystals, Science 361, 582 (2018).
- Brites et al. [2012] C. D. S. Brites, P. P. Lima, N. J. O. Silva, A. Millán, V. S. Amaral, F. Palacio, and L. D. Carlos, Thermometry at the nanoscale, Nanoscale 4, 4799 (2012).
- Quintanilla and Liz-Marzán [2018] M. Quintanilla and L. M. Liz-Marzán, Guiding Rules for Selecting a Nanothermometer, Nano Today 19, 126 (2018).
- Balandin et al. [2008] A. A. Balandin, S. Ghosh, W. Bao, I. Calizo, D. Teweldebrhan, F. Miao, and C. N. Lau, Superior Thermal Conductivity of Single-Layer Graphene, Nano Letters 8, 902 (2008).
- Ghosh et al. [2008] S. Ghosh, I. Calizo, D. Teweldebrhan, E. P. Pokatilov, D. L. Nika, A. A. Balandin, W. Bao, F. Miao, and C. N. Lau, Extremely high thermal conductivity of graphene: Prospects for thermal management applications in nanoelectronic circuits, Applied Physics Letters 92, 151911 (2008).
- Braun et al. [2022] O. Braun, R. Furrer, P. Butti, K. Thodkar, I. Shorubalko, I. Zardo, M. Calame, and M. L. Perrin, Spatially mapping thermal transport in graphene by an opto-thermal method, npj 2D Materials and Applications 6, 1 (2022).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. L. Hollenberg, The nitrogen-vacancy colour centre in diamond, Physics Reports 528, 1 (2013).
- Neumann et al. [2013] P. Neumann, I. Jakobi, F. Dolde, C. Burk, R. Reuter, G. Waldherr, J. Honert, T. Wolf, A. Brunner, J. H. Shim, D. Suter, H. Sumiya, J. Isoya, and J. Wrachtrup, High-precision nanoscale temperature sensing using single defects in diamond, Nano Letters 13, 2738 (2013).
- Toyli et al. [2013] D. M. Toyli, C. F. d. l. Casas, D. J. Christle, V. V. Dobrovitski, and D. D. Awschalom, Fluorescence thermometry enhanced by the quantum coherence of single spins in diamond, Proceedings of the National Academy of Sciences 110, 8417 (2013).
- Kucsko et al. [2013] G. Kucsko, P. C. Maurer, N. Y. Yao, M. Kubo, H. J. Noh, P. K. Lo, H. Park, and M. D. Lukin, Nanometre-scale thermometry in a living cell, Nature 500, 54 (2013).
- Acosta et al. [2013] V. M. Acosta, E. Bauch, M. P. Ledbetter, A. Waxman, L.-S. Bouchard, and D. Budker, Temperature dependence of the nitrogen-vacancy magnetic resonance in diamond, Physical Review Letters 104, 070801 (2013).
- Barry et al. [2020] J. F. Barry, J. M. Schloss, E. Bauch, M. J. Turner, C. A. Hart, L. M. Pham, and R. L. Walsworth, Sensitivity optimization for nv-diamond magnetometry, Rev. Mod. Phys. 92, 015004 (2020).
- Wojciechowski et al. [2018] A. M. Wojciechowski, M. Karadas, C. Osterkamp, S. Jankuhn, J. Meijer, F. Jelezko, A. Huck, and U. L. Andersen, Precision temperature sensing in the presence of magnetic field noise and vice-versa using nitrogen-vacancy centers in diamond, Applied Physics Letters 113, 013502 (2018).
- Gottscholl et al. [2021] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, D. Krauße, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Spin defects in hBN as promising temperature, pressure and magnetic field quantum sensors, Nature Communications 12, 4480 (2021).
- Beaufils et al. [2018] C. Beaufils, W. Redjem, E. Rousseau, V. Jacques, A. Y. Kuznetsov, C. Raynaud, C. Voisin, A. Benali, T. Herzig, S. Pezzagna, J. Meijer, M. Abbarchi, and G. Cassabois, Optical properties of an ensemble of g-centers in silicon, Physical Review B 97, 035303 (2018).
- Kraus et al. [2014] H. Kraus, V. A. Soltamov, F. Fuchs, D. Simin, A. Sperlich, P. G. Baranov, G. V. Astakhov, and V. Dyakonov, Magnetic field and temperature sensing with atomic-scale spin defects in silicon carbide, Scientific Reports 4, 5303 (2014).
- Anisimov et al. [2016] A. N. Anisimov, D. Simin, V. A. Soltamov, S. P. Lebedev, P. G. Baranov, G. V. Astakhov, and V. Dyakonov, Optical thermometry based on level anticrossing in silicon carbide, Scientific Reports 6, 33301 (2016).
- Zhou et al. [2017] Y. Zhou, J. Wang, X. Zhang, K. Li, J. Cai, and W. Gao, Self-protected thermometry with infrared photons and defect spins in silicon carbide, Physical Review Applied 8, 044015 (2017).
- [41] See Supplemental Material at … for further details on experimental and theoretical methods and additional data, which includes Refs. [62-82].
- Zhang et al. [2021] S.-C. Zhang, Y. Dong, B. Du, H.-B. Lin, S. Li, W. Zhu, G.-Z. Wang, X.-D. Chen, G.-C. Guo, and F.-W. Sun, A robust fiber-based quantum thermometer coupled with nitrogen-vacancy centers, Review of Scientific Instruments 92, 044904 (2021).
- Cambria et al. [2023] M. C. Cambria, G. Thiering, A. Norambuena, H. T. Dinani, A. Gardill, I. Kemeny, V. Lordi, A. Gali, J. R. Maze, and S. Kolkowitz, Physically motivated analytical expression for the temperature dependence of the zero-field splitting of the nitrogen-vacancy center in diamond, Phys. Rev. B 108, L180102 (2023).
- Li et al. [2012] W. Li, N. Mingo, L. Lindsay, D. A. Broido, D. A. Stewart, and N. A. Katcho, Thermal conductivity of diamond nanowires from first principles, Physical Review B—Condensed Matter and Materials Physics 85, 195436 (2012).
- Fugallo et al. [2013] G. Fugallo, M. Lazzeri, L. Paulatto, and F. Mauri, Ab initio variational approach for evaluating lattice thermal conductivity, Phys. Rev. B 88, 045430 (2013).
- Swinkels et al. [2015] M. v. Swinkels, M. Van Delft, D. Oliveira, A. Cavalli, I. v. Zardo, R. Van Der Heijden, and E. Bakkers, Diameter dependence of the thermal conductivity of inas nanowires, Nanotechnology 26, 385401 (2015).
- Martin et al. [2009] P. Martin, Z. Aksamija, E. Pop, and U. Ravaioli, Impact of phonon-surface roughness scattering on thermal conductivity of thin si nanowires, Physical review letters 102, 125503 (2009).
- Hochbaum et al. [2008] A. I. Hochbaum, R. Chen, R. D. Delgado, W. Liang, E. C. Garnett, M. Najarian, A. Majumdar, and P. Yang, Enhanced thermoelectric performance of rough silicon nanowires, Nature 451, 163 (2008).
- Li et al. [2003] D. Li, Y. Wu, P. Kim, L. Shi, P. Yang, and A. Majumdar, Thermal conductivity of individual silicon nanowires, Applied Physics Letters 83, 2934 (2003).
- Maire et al. [2017] J. Maire, R. Anufriev, and M. Nomura, Ballistic thermal transport in silicon nanowires, Scientific Reports 7, 41794 (2017).
- Gandolfi et al. [2020] M. Gandolfi, C. Giannetti, and F. Banfi, Temperonic crystal: A superlattice for temperature waves in graphene, Phys. Rev. Lett. 125, 265901 (2020).
- Raya-Moreno et al. [2022a] M. Raya-Moreno, X. Cartoixà, and J. Carrete, BTE-Barna: An extension of almaBTE for thermal simulation of devices based on 2D materials, Computer Physics Communications 281, 108504 (2022a).
- Raya-Moreno et al. [2022b] M. Raya-Moreno, J. Carrete, and X. Cartoixà, Hydrodynamic signatures in thermal transport in devices based on two-dimensional materials: An ab initio study, Physical Review B 106, 014308 (2022b).
- Dragašević and Simoncelli [2024] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors (2024), arXiv:2303.12777 [cond-mat.mtrl-sci] .
- Li et al. [2014] W. Li, J. Carrete, N. A. Katcho, and N. Mingo, Shengbte: A solver of the boltzmann transport equation for phonons, Computer Physics Communications 185, 1747 (2014).
- Ratsifaritana and Klemens [1987] C. Ratsifaritana and P. Klemens, Scattering of phonons by vacancies, International journal of thermophysics 8, 737 (1987).
- Bayn et al. [2011] I. Bayn, A. Bolker, C. Cytermann, B. Meyler, V. Richter, J. Salzman, and R. Kalish, Diamond processing by focused ion beam—surface damage and recovery, Applied Physics Letters 99, 183109 (2011).
- Katcho et al. [2014] N. A. Katcho, J. Carrete, W. Li, and N. Mingo, Effect of nitrogen and vacancy defects on the thermal conductivity of diamond: An ab initio green’s function approach, Phys. Rev. B 90, 094117 (2014).
- Anthony et al. [1991] T. R. Anthony, J. L. Fleischer, J. R. Olson, and D. G. Cahill, The thermal conductivity of isotopically enriched polycrystalline diamond films, Journal of Applied Physics 69, 8122 (1991).
- Tian et al. [2024] S. Tian, T. Wang, H. Chen, D. Ma, and L. Zhang, Phonon coherent transport leads to an anomalous boundary effect on the thermal conductivity of a rough graphene nanoribbon, Phys. Rev. Appl. 21, 064005 (2024).
- Polanco et al. [2023] C. A. Polanco, A. van Roekeghem, B. Brisuda, L. Saminadayar, O. Bourgeois, and N. Mingo, The phonon quantum of thermal conductance: Are simulations and measurements estimating the same quantity?, Science Advances 9, eadi7439 (2023).
- Chambers [1950] R. Chambers, The conductivity of thin wires in a magnetic field, Proceedings of the Royal Society of London. Series A. Mathematical and Physical Sciences 202, 378 (1950).
- Ziman [2001] J. M. Ziman, Electrons and phonons: the theory of transport phenomena in solids (Oxford university press, 2001).
- Mi et al. [2019] S. Mi, A. Toros, T. Graziosi, and N. Quack, Non-contact polishing of single crystal diamond by ion beam etching, Diamond and Related Materials 92, 248 (2019).
- Burek et al. [2014] M. J. Burek, Y. Chu, M. S. Z. Liddy, P. Patel, J. Rochman, S. Meesala, W. Hong, Q. Quan, M. D. Lukin, and M. Lončar, High quality-factor optical nanocavities in bulk single-crystal diamond, Nature Communications 5, 5718 (2014).
- Graziosi et al. [2018] T. Graziosi, S. Mi, M. Kiss, and N. Quack, Single crystal diamond micro-disk resonators by focused ion beam milling, Apl Photonics 3 (2018).
- Riedrich-Möller et al. [2012] J. Riedrich-Möller, L. Kipfstuhl, C. Hepp, E. Neu, C. Pauly, F. Mücklich, A. Baur, M. Wandt, S. Wolff, M. Fischer, S. Gsell, M. Schreck, and C. Becher, One- and two-dimensional photonic crystal microcavities in single crystal diamond, Nature Nanotechnology 7, 69 (2012).
- Cai et al. [2019] Y.-Y. Cai, E. Sung, R. Zhang, L. J. Tauzin, J. G. Liu, B. Ostovar, Y. Zhang, W.-S. Chang, P. Nordlander, and S. Link, Anti-stokes emission from hot carriers in gold nanorods, Nano Letters 19, 1067 (2019), publisher: American Chemical Society.
- Carattino et al. [2018] A. Carattino, M. Caldarola, and M. Orrit, Gold Nanoparticles as Absolute Nanothermometers, Nano Letters 18, 874 (2018).
- Jollans et al. [2020] T. Jollans, M. Caldarola, Y. Sivan, and M. Orrit, Effective Electron Temperature Measurement Using Time-Resolved Anti-Stokes Photoluminescence, The Journal of Physical Chemistry A 124, 6968 (2020).
- Crampton et al. [2018] K. T. Crampton, A. Fast, E. O. Potma, and V. A. Apkarian, Junction Plasmon Driven Population Inversion of Molecular Vibrations: A Picosecond Surface-Enhanced Raman Spectroscopy Study, Nano Letters 18, 5791 (2018).
- Dubi and Sivan [2019] Y. Dubi and Y. Sivan, “hot” electrons in metallic nanostructures—non-thermal carriers or heating?, Light: Science & Applications 8, 89 (2019).
- Ziegler [1985] J. P. Ziegler, James F.and Biersack, The stopping and range of ions in matter, in Treatise on Heavy-Ion Science, Vol. 6, edited by D. A. Bromley (Springer US, Boston, MA, 1985) pp. 93–129.
- Peierls [1929] R. Peierls, Zur kinetischen theorie der wärmeleitung in kristallen, Annalen der Physik 395, 1055 (1929).
- Peierls [1996] R. E. Peierls, Quantum theory of solids (Clarendon Press, 1996).
- Simoncelli et al. [2019] M. Simoncelli, N. Marzari, and F. Mauri, Unified theory of thermal transport in crystals and glasses, Nature Physics 15, 809 (2019).
- Simoncelli et al. [2022] M. Simoncelli, N. Marzari, and F. Mauri, Wigner formulation of thermal transport in solids, Physical Review X 12, 041011 (2022).
- Di Lucente et al. [2023] E. Di Lucente, M. Simoncelli, and N. Marzari, Crossover from boltzmann to wigner thermal transport in thermoelectric skutterudites, Physical Review Research 5, 033125 (2023).
- Cepellotti and Marzari [2016] A. Cepellotti and N. Marzari, Thermal transport in crystals as a kinetic theory of relaxons, Physical Review X 6, 041013 (2016).
- Cepellotti et al. [2015] A. Cepellotti, G. Fugallo, L. Paulatto, M. Lazzeri, F. Mauri, and N. Marzari, Phonon hydrodynamics in two-dimensional materials, Nature communications 6, 6400 (2015).
- Note [1] In particular, Eq. 6 of Ref. [44] encompasses as special cases both thin films and thin wires, as discussed after Eq. 5 of Ref. [62] and Eq. 11.2.9 of Ref. [63].
- Mathematica [2019] Mathematica, Wolfram Research Inc., Champaign, IL (Version 12.0, 2019).
I Supplementary materials
I.1 Sample preparation
We use commercial (100) CVD single crystalline diamond substrate (Hiqute Diamond, Rouge-T12-100), enriched to above 99.95% in 12C isotopes and containing an NV-rich layer (1–2 ppm) of µm thickness at the surface. We first perform preliminary acid cleaning: piranha solution (H2SO4:H2O2 3:1) and hydrofluoric acid remove organics and metallic residuals, respectively. The substrate is then mounted on a silicon chip with 2 µm thick wet oxide via QuickStick wax. We apply a reported non-contact ion beam polishing method to improve the surface roughness [64].
To define the chromium patches, a 150 nm-thick Cr layer is evaporated on the NV-rich diamond surface. The patches are then defined using electron beam lithography followed by wet etching. 150 nm of hydrogen silsesquioxane (HSQ; XR-1541-006) is spin-coated, written with electron beam lithography, and subsequently developed in tetramethylammonium hydroxide (TMAH 25) as a mask. TechniEtch Cr01 from MicroChemicals ((NH4)2Ce(NO3)6+HClO4) removes the unmasked Cr layer. Non-contact ion polishing is used to clean the sample surface, and then the patches are released by HF 1%.
We then define the cantilevers, aligned to the Cr patches, using a second round of electron beam lithography followed by several etching steps. A thin titanium layer (40 nm) is evaporated on the surface. A 2 µm-thick layer of HSQ (FOx-16) is spin coated above and exposed with electron beam lithography followed by the aforementioned development process. The Ti layer serves as a conductive and adhesive layer between the diamond the HSQ. A Chlorine-based reactive ion etching process allows to pattern the Ti layer (STS Multiplex ICP, 800 W ICP power, 150 W bias power, 10 sccm Cl2 and 10 sccm BCl3, 3 mTorr). To transfer the pattern from the HSQ hard mask to the diamond, a vertical etch is applied with directional oxygen plasma (STS Multiplex ICP, 400 W ICP power, 200 W bias power, 50 sccm O2, 15 mTorr). Faraday cage angled etch is then used under the same plasma condition as the previous step to undercut the structures [65]. The hard mask is stripped in the HF 1% bath.
Since the Faraday cage angled etch technique restricts the lateral dimensions of suspended structures, we proceed further with focused ion beam (FIB) milling. This method allows greater flexibility in structuring, enabling us to precisely polish narrow cantilevers and undercut wider ones [66, 67]. A 200 nm Ti layer is deposited to shield the sample from gallium (Ga) implantation and to prevent drift during milling caused by charge accumulation. The milling is conducted with a Zeiss CrossBeam 540 SEM/FIB dual beams system using 30 kV Ga+ ions on both sides of each cantilever. We apply cross-section mode with 1.5 nA beam current to sputter away most of the material and lower the ion current to 700 pA to realize better sidewall smoothness. Finally, cantilevers with finely polished sidewalls are revealed after removing the Ti layer in HF bath, as shown in Fig. 1(b).
I.2 Experimental setup
All experiments are carried out in our home-built confocal microscope. The diamond sample is mounted on a temperature-controlled stage, kept at a constant temperature of 22 for experiments in Fig. 2 and Fig. 4. The output beam of a 515nm laser is split into two beams, one for heating the cantilever and the other for readout of the NV centers’ ZFS. The power of each beam can be controlled independently. The two beams are focused on different positions of the cantilevers with a high numerical aperture microscope objective (NA=0.9). PL of the NV centers is collected into a multi-mode fiber and detected with a multi-pixel photon counter (Hamamatsu C14455-1550GA). Microwave are delivered to the sample with a wire loop placed in close vicinity to the diamond sample. Microwave frequency is modulated directly within the microwave generator (Rhodes&Schwarz SMB 100A), with sinusoidal modulation shape, to use in combination with the lock-in amplifier. The modulation period is 0.2 ms.
The readout parameters are chosen to optimize temperature sensitivity: the readout laser power is 0.2 mW, microwave power 41 dBm and microwave frequency modulation depth 5 MHz (corresponding roughly to half the ODMR linewidth). We have checked that the above parameters do not affect the readout ZFS value, confirming that the readout does not induce local heating of the cantilever. We have also verified that the heating laser does not induce any change in ZFS when focused on other portions of the cantilever, away from the metal patch.
I.3 Independent study of the patch temperature
To confirm that the metal patch is well thermalized to the diamond cantilever, we measure the hot carrier anti-Stokes emission of the patch to get an independent estimate of its temperature. Luminescence in noble metals following photoexcitation is a complex process influenced by various mechanisms, including electron-electron scattering, electron-phonon coupling, multiphoton excitations, and surface plasmons[68, 69]. Anti-Stokes emission corresponds to photons emitted at a higher energy than the energy of the laser incident on the material. As detailed in previous studies, the anti-Stokes emission spectrum can be fit with Boltzmann statistics to determine the effective temperature [68, 70]. The spectrum is given by:
where and represent emission probability and a density of states respectively and represents the anti-Stokes temperature. The exponential term, or Boltzmann factor, merges the energy contribution from electrons and phonons. The term varies much slower than the Boltzmann factor [71, 70], hence Eq. (S1) can be simplified as:
from which we can determine the effective temperature .
We measure the anti-Stokes emission spectrum on three different cantilevers, using 532nm continuous wave excitation. This experiment was performed on a different sample than discussed in the main text, where the metal patch was made of gold (Au) instead of Cr. Still, a few nm-thin Cr layer was used as adhesion layer between Au and diamond. Hence, aside from the absorption coefficient of the patch , this difference is not expected to lead to any major qualitative difference in local heat generation and patch thermalization. Anti-stokes emission is measured on a spectrometer. A notch filter rejects most of the reflected excitation laser. Collected spectra are shown in Fig. S1 (a) and can be well fitted with the exponential decay from Eq. (S2) over the range that we can experimentally access. Note that we discard the energy window corresponding to the Raman peak of diamond, to avoid spurious residual contribution from the diamond cantilever. We extract the anti-Stokes temperature from the fit.
The relationship between and the lattice temperature is in general not straightforward, specifically under continuous wave illumination [72]. Hence, we perform a calibration measurement to relate the extracted to the patch temperature, using the temperature-controlled stage on which the diamond sample is mounted (see inset of Fig. S1(b)). We find that the relation between and the controlled patch temperature can be well described using a linear fit: . We can therefore determine the temperature of gold patches versus heating laser power, on cantilevers with different widths, as shown in Fig. S1(b). Although this method does not guarantee high precision (note also that the extracted zero-laser-power temperature is significantly lower on the 1.18 µm cantilever), this measurement proves that the metal patch temperature increases linearly when heating laser is applied. Temperature estimates with anti-Stokes emission are also found to be on the same order as values presented in the main text on diamond cantilevers for similar heating power, meaning that the interface thermal resistance between the patch and diamond is small. Consequently, convective and radiative thermal losses in the patch can safely be neglected.
I.4 Estimation of experimental effective thermal conductivity
The effective thermal conductivity of our cantilevers is defined as . The temperature gradient is directly extracted from the measured temperature profile. We must only estimate the heat flux density in order to obtain . As introduced in the main text, we have . The heating laser power incident on the Cr patch is measured directly. The Cr patch absorption coefficient is estimated by measuring its reflection coefficient and assuming negligible transmission and scattering, i.e. . Finally, the cantilever cross-section is , where both the width and cross-section angle are estimated from SEM images of the different cantilevers, see for example Fig. 3(a) of the main text.
Note that we use an uncertainty of 50 nm in the extracted width from SEM images. This uncertainty needs to be properly taken into account in the uncertainty of , together with the uncertainty on the temperature gradient. We find that the width uncertainty is the dominant factor, explaining the significantly bigger error bars in , see Fig. 4, compared to the temperature gradient in Fig. 3(c).
I.5 Additional figures
I.6 Theoretical analysis
An established approach to describe thermal transport in crystals from first principles is the linearized phonon Boltzmann transport equation (LBTE) [74, 75, 63]:
where is phonon distribution (, with being the phonon wavevector and the phonon mode index). The LBTE is a microscopic model, which describes the propagation and scattering of phonons in the presence of a temperature gradient. The LBTE can be parametrized from first principles, and its linear-response solution allows to predict the thermal conductivity. We note, in passing, that the LBTE has been recently generalized to account not only for the ability of phonons to propagate particle-like, but also also for their capability to tunnel wave-like [76, 77, 78]. However, in diamond the LBTE is sufficient to accurately describe thermal transport, since in this case propagation dominates over tunneling [12].
The first-principles solution of the LBTE has the drawback of being computationally expensive, making it unpractical to explore how shape effects affect heat transport. Therefore, to investigate thermal transport in cantilevers having different shapes and sizes, we rely on the mesoscopic viscous heat equations (VHE) [12]. The VHE are two coupled partial differential equations for the mesoscopic temperature () and phonon drift-velocity () fields. They are obtained from a coarse-graining the LBTE[79, 12]. The VTE encompass both the diffusive (Fourier) and hydrodynamic regimes [80, 79] of thermal transport, as well as the crossover between them. It has been shown that the VHE rationalizes the experimental demonstration of hydrodynamic transport in graphite [12], and also yields predictions compatible with the space-dependent solution of the LBTE[52, 53] in complex-shaped devices [54]. In the steady-state regime, for a material with isotropic transport properties such as diamond, the VHE read:
where is the phonon viscosity, the coupling coefficients and arise from the energy-crystal momentum relation for phonons [12], while accounts for the presence of momentum-dissipating scattering processes [12]. We use capital to emphasize the space variation at the mesoscopic scale.
From Eq. (S4) one can derive energy and momentum balance equations in terms of a heat flux generalized to the hydrodynamic regime:
where labels Cartesian directions.
The mesoscopic heat flux is then used to evaluate the effective thermal conductivity across the cantilevers. For a cantilever, taken along axis, of length and with an imposed temperature difference at its boundaries , we get:
where is the the average heat flux discussed in the main text.
In the limit of negligible hydrodynamic effects, Eqs. (S4) reduce to the Fourier’s law of diffusive heat conduction:
In this regime the heat flux simply becomes
The importance on viscuous effects can thus be estimated by comparing the results of computing the heat flux with Eq. (S5) (full solution) versus Eq. (S8) (purely diffusive case).
All the parameters entering the VHE can be determined from the first-principles solution of the full LBTE (i.e. accounting for the actual phonon band structure and full collision matrix). An exact solution of the LBTE has been discussed for the case of nanowires [44, 55] under the assumption of diffusive boundaries. This approach allows to account for the influence of boundary scattering on the conductivity of wires, providing an expression which depends only on the cross-section of the wire 111In particular, Eq. 6 of Ref. [44] encompasses as special cases both thin films and thin wires, as discussed after Eq. 5 of Ref. [62] and Eq. 11.2.9 of Ref. [63] and can be applied to the triangular cross-section nanowire studied in this work. After verifying with the established Casimir model [45] that boundary scattering has strong effects on the conductivity but has negligible effect on the other parameters entering the VHE (including the thermal viscosity ), we accounted for nanowire boundary scattering on the conductivity only and used bulk values (calculated in Ref. [12]) for the other parameters entering in the VHE. This approximation is further justified a posteriori, since Fig. 4 shows that viscous effects are unimportant in the cases considered here.
Finally, we discuss how we considered the Ga contamination in the vicinity of milled faces of the cantilevers (i.e. bottom faces, see inset of Fig. 4). Since with the methods currently available it is not possible to solve the LBTE with position-dependent Ga content, we approximatively consider uniform Ga impurities over the whole cantilever cross-section. This approximation allows to capture the increased influence of the Ga-contaminated layer (of fixed depth , see main text) when reducing the cantilever width . In particular, we consider an impurity concentration proportional to the ratio of contaminated area to total cross-section area, referred to as penetration ratio
where the approximated equality holds at linear order in ; details on the geometrical analysis performed to compute PR are reported in the next section. Eq. (S9) allows us to describe the width-dependent average Ga concentration as:
where is the concentration of Ga impurities in the contaminated layer. We thus determine the conductivity from the linear-response solution of the LBTE, accounting for phonon-impurity scattering with impurity strength parameter corresponding to Ga impurities with concentration [56].
In summary, we determine the linear-response solution of the LBTE (Eq. (S3)) for diamond nanowires with width-dependent Ga impurity scattering using the ShengBTE package [55] and use the bulk thermal viscosity computed following the theoretical prescriptions of Ref. [12]. Then, we solve the VHE numerically (Eq. (S4)) using a finite-element solver implemented in Mathematica [82], imposing a temperature gradient K around room temperature on µm long cantilevers. Finally, we obtain the effective thermal conductivity from Eqs (S5) and (S6). The Ga concentration in contaminated areas is set in the ShengBTE simulation as to fit the experimental trend of vs width (black squares in Fig. 4 of the main text). This is the sole fitting parameter in our model.
I.7 Evaluation of the Ga participation ratio
Given that Ga impurities are mostly localized on the two slanted sides of Diamond cantilevers (see inset of Fig. 4), we can compute the participation ratio, PR, as the ratio between the azure area in the triangle on the right of Fig. 4, , and the total triangular section, . Considering an angle at the surface equal to , with some algebra one finds
where the fraction multiplied by a factor of 2 represents each of the -deep areas related to the two slanted sides (see the inset of Fig. 4 of the main text). On the other hand, the total triangular section is simply
These considerations allow us to approximate to linear order in and justify Eq. (S9) as shown below:
The solution of the LBTE with a space-dependent scattering term for nanowires is obtained using the ShengBTE package. Even if the solution is valid for any cross-section shape (it depends only on the cross-section value), the ShengBTE implementation is coded assuming a circular cross-section. So, the widths of the prismatic cantilevers are obtained from the ShengBTE simulations by matching circular cross-section used in the simulation, ( being the radius of the cylindrical nanowire simulated), with the triangular cross-section, . This simply translates into getting the width values of triangular nanowires (used on the x axis of Fig. 4 of the main text) as .
Finally, in Fig. S7, we show further effective thermal conductivity calculations obtained using a fixed (non width-resolved) Ga-impurity concentration. We clearly observe that the slope of such curves does not capture the experimental trend, highlighting the necessity to consider a Ga-impurity model.