### v3

# Imaging heat transport in suspended diamond nanostructures with integrated spin defect thermometers

Abstract
Among all materials, mono-crystalline diamond has one of the highest measured thermal conductivities, with values above 2000 W/m/K at room temperature. This stems from momentum-conserving ‘normal’ phonon-phonon scattering processes dominating over momentum-dissipating ‘Umklapp’ processes, a feature that also suggests diamond as an ideal platform to experimentally investigate phonon heat transport phenomena that violate Fourier’s law. Here, we introduce dilute nitrogen-vacancy color centers as in-situ, highly precise spin defect thermometers to image temperature inhomogeneities in single-crystal diamond microstructures heated from ambient conditions. We analyze cantilevers with cross-sections in the range from about 0.2 to 2.6 µm2, observing a strong reduction of the cantilevers’ conductivity as width decreases. We use first-principles simulations based on the linearized phonon Boltzmann transport equation and viscous heat equations to quantitatively predict the cantilevers’ thermal transport properties, rationalizing how the interplay between intrinsic and extrinsic phonon scattering mechanisms determines the observed non-diffusive behavior. Our temperature-imaging method paves the way for the exploration of unconventional, non-diffusive heat transport phenomena in devices and nanostructures of arbitrary geometries.
Carbon-based materials such as graphene, graphite or diamond present exceptional heat-conduction properties. Due to large phonon group velocities, they exhibit high thermal conductivity – with diamond reaching record values at room temperature and below [1, 2, 3]. Moreover, the prevalence of normal (momentum-conserving) phonon-phonon scattering processes over Umklapp (momentum non-conserving) processes in these materials leads to phonon heat transport phenomena that violate Fourier’s law for macroscopic diffusive heat propagation [4]. In graphene and graphite, ballistic transport in nanoscale devices [5, 6] or hydrodynamic heat transport [7, 8, 9, 10, 11, 12, 13] have been evidenced at or near room temperature. However, theoretical description of non-diffusive transport phenomena when both ballistic (extrinsic) and hydrodynamic (intrinsic) effects compete remains challenging, as it relies on phenomenological models that do not have a rigorous first-principles grounding [14, 15, 13]. In diamond, the influence of phonon scattering at grain boundaries has been explored in polycrystalline samples [16] and unexpected size effects were recently reported in the thermal conductivity of microparticles [17], but experimental data on low-dimensional single-crystal diamond is lacking. Understanding heat transport in nanostructured diamond will be valuable in view of its applications in optomechanics [18, 19], nonlinear optics [20], quantum nanophotonics [21], as well as in heat management devices [22].
A major challenge in the study of thermal transport at the nanoscale is to find appropriate temperature sensors. Traditional measurements of the heat conductivity rely on resistive sensors combined with advanced sample engineering [5, 11] and cannot provide spatially-resolved temperature maps. Scanning thermal microscopy offers excellent spatial resolution, but is invasive and requires challenging calibration of the probe-sample contact thermal resistance [23, 24]. Non contact optical thermometry techniques such as transient thermal gratings [25, 9], and time- or frequency-domain thermoreflectance [26, 27, 28] have been widely employed on simple planar geometries but require the deposition of a transducer material, typically a metal [29], and don’t provide a direct temperature readout [30]. Other optical techniques [31, 32], including Raman sideband thermometry [33, 34, 35, 36], offer good spatial resolution but rather poor temperature accuracy, typically not better than a few Kelvin.
In this work, we use the spin resonance of dilute color centers in the diamond lattice as precise in-situ temperature sensors to investigate heat transport properties of suspended diamond microstructures. We then obtain quantitative agreement with theoretical predictions from multiscale first-principle modeling. Diamond hosts optically active spin defects, among which the negatively charged nitrogen vacancy (NV) center [37], which has been used as in-situ nanoscale thermometer [38, 39, 40], also in combination with a scanning tip [41]. Negatively charged NV centers are spin-1 atomic defects formed by a substitutional nitrogen atom neighboring a vacancy in the diamond lattice. Their spin can be initialized and readout by optical pumping and monitoring the photoluminescence intensity. The resonance frequency associated with the transition between and spin states is temperature-dependent [42], as can be tracked with either continuous-wave or pulsed optically detected magnetic resonance (ODMR) [43]. NV centers are thus excellent thermometers, offering nanoscale spatial resolution and sub- sensitivity [44] for steady-state temperature measurement near room temperature. Solid-state spin defects acting as temperature sensors are also found in a broad range of materials, including hBN [45], Si [46] and SiC [47, 48, 49], making this method widely applicable.
The principle of our experiment is summarized in Fig.1a. We fabricate suspended micro-cantilevers out of single crystal diamond grown by chemical vapour deposition and enriched in (%), containing NV centers up to a concentration of 3 ppm (HiQuTe Diamond). Cantilevers are patterned using electron-beam lithography and subsequent etching, in two steps: first a step of vertical, oxygen-based reactive ion etching, followed by focused ion beam (FIB) milling at an angle from the diamond surface. This results in triangular cantilever cross-sections, as visible in Fig. 1b. The cantilever acts as a one-dimensional channel for heat transport. A small chromium patch is added at the cantilever tip to serve as a heat source upon absorption of a 515 nm laser focused onto it (referred to as heating laser) and the bulk diamond acts as a heat sink. Complete details on the sample fabrication can be found in the Supplementary Material (SM) [50].
Photoluminescence from NV centers is collected in confocal geometry, using a second 515 nm laser beam to address a large ensemble of NV centers (about ) located in the diffraction-limited spot. This second laser is referred to as readout laser. The temperature-dependent zero-field splitting (ZFS) between and states is measured through continuous wave (cw) ODMR. We use a lock-in amplifier (LIA) combined with frequency modulation of the microwave to achieve high signal to noise ratio [51]. We first calibrate the variation of ZFS, , against a reference bath temperature applied to the entire sample. Fig. 1c presents measured both in bulk diamond and on cantilevers. Even though the absolute value of is found to depend on position [50], possibly due to the presence of local strain, is position-independent. Indeed, since our cantilevers are free-standing, we do not expect that heating induces any additional strain (except, maybe, very near its base). Our data is fitted by the model from Ref. [52], with only two fitting parameters (dashed line in Fig. 1c).
The calibration is used to convert values of to local temperature under the readout laser spot. Noise analysis gives a typical temperature sensitivity of 0.27 for the temperature readout on the cantilever [50]. The choice of cw-ODMR is a trade-off between sensitivity requirement and simplicity of experimental implementation. Even though more sophisticated NV-based thermometry schemes allow to reach better sensitivity, relying e.g. on pulse sequences [38, 39] or on multi-frequency drive [44], the dominant sources of uncertainties in our case come from the estimates of absorbed laser power and cantilever cross-section, as detailed below.
The heating laser is used to induce a heat flow in a cantilever. Fig. 2a presents the measured local temperature rise as a function of heating laser power , for a readout spot located at a fixed position in the middle of the cantilever (see Fig. 3a,II). We observe a linear relation between and heating power. Standard deviations of residuals after a linear fit give a temperature error of K [50]. This value includes systematic errors due e.g to imperfect alignment of the heating laser on the chromium patch, explaining why it is slightly higher than the readout error ( K with our 3 s integration time).
Fixing the heating laser power, we measure the space-dependent temperature profile along the cantilever by sweeping the position of the readout laser. For each readout laser position, the ZFS values are measured without and with heating. To illustrate the method, the temperature profile for a cantilever of varying width is shown in Fig. 2c. Each section of constant width shows a linear temperature profile. The narrow section (from 4 to 10 µm) has a much steeper temperature gradient than the wide section (0 to 4 µm), while the temperature is constant on the free-hanging side ( to 0 µm). In our experimental setting, the spatial resolution is diffraction-limited, the readout laser spot size being 0.44 µm in diameter.
To better understand this temperature profile, we consider the expected behavior for standard diffusive heat transport. According to Fourier’s law, in 1D, , with the local heat flux density and the material’s thermal conductivity, which depends slowly on temperature. We confirmed that the Cr patch’s temperature remains close to that of the underlying diamond [50], so that convective and radiative thermal losses can be neglected. In this case, , with the heating laser power incident on the Cr patch, the Cr patch absorption coefficient and the cantilever cross-section. For small , Fourier’s law predicts a linear temperature profile along the cantilever, compatible with the results from Fig. 2. However, having a linear temperature profile is a necessary but not sufficient condition for the validity of Fourier’s law, since it also assumes thermal conductivity to be an intrinsic, size-independent material property, and consequently predicts a heat flux that scales linearly with the cantilever cross-section. Instead, in diamond, has been predicted to decrease to approximately half of its bulk value at widths 0.7m [53, 54] – a much stronger reduction than other materials for this channel size[55, 56, 57, 58, 59]. We now show that a reduction of is observed when reducing the lateral size of the cantilever.
We investigate cantilevers with different lateral width , in the range 0.8-2.5 µm, and identical length of 10 µm. The lower bound of this range originates from limitations in temperature sensitivity for cantilever widths below the diffraction limited spot size (due to the reduced number of probed NV centers). Conversely, the upper bound is due to the difficulty in fabricating suspended cantilevers with dimensions bigger than 3 µm (due to the long etching time required). We measure the temperature profile along each individual cantilever, adjusting such that the maximum temperature rise is comparable in all cases and remains below 40 K, within the calibration range of Fig. 1. The results are presented in Fig. 3, with panel (c) presenting the measured temperature gradient , normalized by , versus cantilever width . In the case of purely diffusive thermal transport, i.e., size-independent , Fourier’s law predicts that this normalized temperature gradient is inversely proportional to the cantilevers cross-section. Specifically, since all our cantilevers have triangular cross-section with same angle , Fourier’s law predicts . In contrast, our experiments show a clear deviation from this prediction for µm. On the other hand, in the limit purely dominated by boundary scattering, thermal conductivity is expected to scale as [60], corresponding to . Our experiments are between these two regimes, indicating the necessity of accounting for both extrinsic (boundary) and intrinsic (normal and Umklapp) phonon scattering mechanisms.
We model thermal transport in our cantilevers using the viscous heat equations (VHE), two coupled mesoscopic heat-transport equations for temperature and phonon drift velocity [61, 62] that generalize Fourier’s laws encompassing not only heat diffusion but also fluid-like behavior [9, 10, 11, 12, 63]. The VHE can be parametrized from first principles and allow to describe how micro-structured geometries influence thermal transport with accuracy comparable to microscopic models based on the full space-dependent solution of the linearized Boltzmann transport equation (LBTE) [64, 65], at much reduced computational cost [62]. Therefore, we employ the VHE to investigate how the geometry of our cantilevers affects thermal transport, and we assess the relevance of viscous and finite-size effects by comparing the VHE predictions with those obtained from Fourier’s law.
We simulate the three-dimensional experimental geometry, imposing a temperature gradient K around room temperature on µm long cantilevers. Parameters entering in the VHE or Fourier’s law are computed from first-principles, accounting for third-order anharmonic phonon-phonon scattering [66, 54], phonon-impurity scattering [67], as well as for the effects of rough cantilever’s boundaries [53] (see SM [50] for details). Importantly, the description of finite-size effects is based on the characteristic length scale of the problem, following Casimir’s treatment of boundary scattering in small devices at low temperatures [68]111We note that this is formally analogous to Bosanquet-type regularizations for the transport coefficients of rarefied fluids [75, 62].. In this regime, where boundary scattering occurs at a rate comparable to or higher than intrinsic anharmonic phonon scattering, the appropriate length scale to use is one-half of the smallest system dimension [70]. Then, we solve the VHE or Fourier’s law and determine the heat flux from their solution. Finally, we evaluate the effective thermal conductivity as , where is the cantilever cross section, and the oriented surface element.
In Fig. 4 we compare the experimental values for , extracted from the measurements in Fig. 3 (see SM [50] for details), against predictions obtained solving the VHE. We obtain good quantitative agreement between simulations and experiments. The striking deviation from diffusion, already discussed in Fig. 3, is visible as a sharp reduction of away from the bulk value in Fig. 4. A more detailed analysis of the strength of viscous hydrodynamic and ballistic effects reveals that the latter are the main contribution to the reduction of (see SM [50]). Heat transport in our cantilever is thus within the “apparent diffusion” regime [71], in the sense that it is reasonably well described by a diffusive (Fourier-like) equation with a non-intrinsic thermal conductivity (i.e., depending on extrinsic factors such as the sample size).
Fig. 4 also includes experimental data for cantilevers fabricated on distinct diamond samples, isotopically enriched in 12C and with the same NV concentration, but without relying on FIB (see SM [50]). All datasets follow the same trend, confirming in particular that possible damage induced by FIB at the cantilevers’ surface plays a negligible role in their transport properties.
In conclusion, we have demonstrated that nearly one-dimensional diamond cantilevers with sub-micrometer widths and embedded NV centers exhibit clear signatures of non-diffusive thermal transport, manifested as an effective thermal conductivity with sub-ballistic size dependence. We have quantitatively explained these observations from a first-principles multiscale approach combining the linear-response solution of the Boltzmann transport equation and the mesoscopic viscous heat equations: sub-ballistic transport arises from the interplay between intrinsic phonon interactions—including both crystal-momentum-conserving“hydrodynamic” scattering events as well as crystal-momentum-dissipating Umklapp scattering events—and extrinsic effects determined by size and geometry. This experiment sets the stage for investigating non-diffusive heat transport phenomena where intrinsic and extrinsic mechanisms coexist. First potential examples are nanostructured geometries that promote the formation of heat vortices [62], or induce thermal rectification [30]. A second example is investigating long-standing questions on the influence of sidewall roughness on phonon-boundary scattering [72, 73], since by varying the etching method and chemical termination of the diamond surface one can engineer these extrinsic parameters. Third, this experimental setup can be pushed toward even smaller channels to address open questions in quantized phononic heat conductance [74]. Overall, our results rationalize non-diffusive, sub-ballistic heat transport in diamond, and more generally demonstrate the potential of solid state spin defects in ultrahigh thermal conductivity materials to explore heat transport phenomena in non-diffusive regimes.
Acknowledgements.
Acknowledgments. This project has received funding from the Swiss National Science Foundation through grant 198898 and 216406, EPFL Center for Quantum Science and Engineering through a Collaborative Research Fellowship, the European Union’s Horizon 2020 research and innovation programmes under the Marie Skłodowska-Curie grant agreement 945363 and under grant agreement 101112347, Project NerveRepack. M. S. acknowledges support from: (i) the Kelvin2 HPC platform at the NI-HPC Centre (funded by EPSRC and jointly managed by Queen’s University Belfast and Ulster University); (ii) the UK National Supercomputing Service ARCHER2, for which access was obtained via the UKCP consortium and funded by EPSRC [EP/X035891/1]; (iii) Gonville and Caius College. N.M and E.D. acknowledge support from the Swiss National Science Foundation (SNSF), through Grant No. CRSII5_189924 (“Hydronics” project). The diamond cantilevers were fabricated at the EPFL Center of MicroNanoTechnology (CMi) and we gratefully acknowledge the support of EPFL CMi.References
- Berman et al. [1975] R. Berman, P. R. W. Hudson, and M. Martinez, Nitrogen in diamond: evidence from thermal conductivity, Journal of Physics C: Solid State Physics 8, L430 (1975).
- Olson et al. [1993] J. R. Olson, R. O. Pohl, J. W. Vandersande, A. Zoltan, T. R. Anthony, and W. F. Banholzer, Thermal conductivity of diamond between 170 and 1200 k and the isotope effect, Phys. Rev. B 47, 14850 (1993).
- Inyushkin et al. [2018] A. V. Inyushkin, A. N. Taldenkov, V. G. Ralchenko, A. P. Bolshakov, A. V. Koliadin, and A. N. Katrusha, Thermal conductivity of high purity synthetic single crystal diamonds, Phys. Rev. B 97, 144305 (2018).
- Chen [2021] G. Chen, Non-fourier phonon heat conduction at the microscale and nanoscale, Nature Reviews Physics , 1 (2021).
- Bae et al. [2013] M.-H. Bae, Z. Li, Z. Aksamija, P. N. Martin, F. Xiong, Z.-Y. Ong, I. Knezevic, and E. Pop, Ballistic to diffusive crossover of heat flow in graphene ribbons, Nature Communications 4, 1734 (2013).
- Nika and Balandin [2012] D. L. Nika and A. A. Balandin, Two-dimensional phonon transport in graphene, Journal of Physics: Condensed Matter 24, 233203 (2012).
- Cepellotti et al. [2015] A. Cepellotti, G. Fugallo, L. Paulatto, M. Lazzeri, F. Mauri, and N. Marzari, Phonon hydrodynamics in two-dimensional materials, Nature communications 6, 6400 (2015).
- Lee et al. [2015] S. Lee, D. Broido, K. Esfarjani, and G. Chen, Hydrodynamic phonon transport in suspended graphene, Nature Communications 6, 6290 (2015).
- Huberman et al. [2019] S. Huberman, R. A. Duncan, K. Chen, B. Song, V. Chiloyan, Z. Ding, A. A. Maznev, G. Chen, and K. A. Nelson, Observation of second sound in graphite at temperatures above 100 k, Science 364, 375 (2019).
- Ding et al. [2022] Z. Ding, K. Chen, B. Song, J. Shin, A. A. Maznev, K. A. Nelson, and G. Chen, Observation of second sound in graphite over 200 k, Nature Communications 13, 285 (2022).
- Machida et al. [2020] Y. Machida, N. Matsumoto, T. Isono, and K. Behnia, Phonon hydrodynamics and ultrahigh–room-temperature thermal conductivity in thin graphite, Science 367, 309 (2020).
- Jeong et al. [2021] J. Jeong, X. Li, S. Lee, L. Shi, and Y. Wang, Transient hydrodynamic lattice cooling by picosecond laser irradiation of graphite, Phys. Rev. Lett. 127, 085901 (2021).
- Huang et al. [2023] X. Huang, Y. Guo, Y. Wu, S. Masubuchi, K. Watanabe, T. Taniguchi, Z. Zhang, S. Volz, T. Machida, and M. Nomura, Observation of phonon poiseuille flow in isotopically purified graphite ribbons, Nature Communications 14, 2044 (2023).
- Aharon-Steinberg et al. [2022] A. Aharon-Steinberg, T. Völkl, A. Kaplan, A. K. Pariari, I. Roy, T. Holder, Y. Wolf, A. Y. Meltzer, Y. Myasoedov, M. E. Huber, et al., Direct observation of vortices in an electron fluid, Nature 607, 74 (2022).
- Palm et al. [2024] M. L. Palm, C. Ding, W. S. Huxter, T. Taniguchi, K. Watanabe, and C. L. Degen, Observation of current whirlpools in graphene at room temperature, Science 384, 465 (2024).
- Sood et al. [2018] A. Sood, R. Cheaito, T. Bai, H. Kwon, Y. Wang, C. Li, L. Yates, T. Bougher, S. Graham, M. Asheghi, M. Goorsky, and K. E. Goodson, Direct Visualization of Thermal Conductivity Suppression Due to Enhanced Phonon Scattering Near Individual Grain Boundaries, Nano Letters 18, 3466 (2018).
- Wang and Sun [2024] Y. Wang and B. Sun, Anomalously strong size effect on thermal conductivity of diamond microparticles, Applied Physics Letters 125, 042202 (2024).
- Tao et al. [2014] Y. Tao, J. M. Boss, B. A. Moores, and C. L. Degen, Single-crystal diamond nanomechanical resonators with quality factors exceeding one million, Nature Communications 5, 3638 (2014).
- Mitchell et al. [2016] M. Mitchell, B. Khanaliloo, D. P. Lake, T. Masuda, J. P. Hadden, and P. E. Barclay, Single-crystal diamond low-dissipation cavity optomechanics, Optica 3, 963 (2016).
- Hausmann et al. [2014] B. J. M. Hausmann, I. Bulu, V. Venkataraman, P. Deotare, and M. Lončar, Diamond nonlinear photonics, Nature Photon 8, 369 (2014).
- Moody et al. [2022] G. Moody, V. J. Sorger, D. J. Blumenthal, P. W. Juodawlkis, W. Loh, C. Sorace-Agaskar, A. E. Jones, K. C. Balram, J. C. F. Matthews, A. Laing, M. Davanco, L. Chang, J. E. Bowers, N. Quack, C. Galland, I. Aharonovich, M. A. Wolff, C. Schuck, N. Sinclair, M. Lončar, T. Komljenovic, D. Weld, S. Mookherjea, S. Buckley, M. Radulaski, S. Reitzenstein, B. Pingault, B. Machielse, D. Mukhopadhyay, A. Akimov, A. Zheltikov, G. S. Agarwal, K. Srinivasan, J. Lu, H. X. Tang, W. Jiang, T. P. McKenna, A. H. Safavi-Naeini, S. Steinhauer, A. W. Elshaari, V. Zwiller, P. S. Davids, N. Martinez, M. Gehl, J. Chiaverini, K. K. Mehta, J. Romero, N. B. Lingaraju, A. M. Weiner, D. Peace, R. Cernansky, M. Lobino, E. Diamanti, L. T. Vidarte, and R. M. Camacho, 2022 roadmap on integrated quantum photonics, Journal of Physics: Photonics 4, 012501 (2022).
- Liu et al. [2015] X. Liu, G. Zhang, and Y.-W. Zhang, Surface-engineered nanoscale diamond films enable remarkable enhancement in thermal conductivity and anisotropy, Carbon 94, 760 (2015).
- Gomès et al. [2015] S. Gomès, A. Assy, and P.-O. Chapuis, Scanning thermal microscopy: A review, physica status solidi (a) 212, 477 (2015).
- Zhang et al. [2020] Y. Zhang, W. Zhu, F. Hui, M. Lanza, T. Borca-Tasciuc, and M. Muñoz Rojo, A review on principles and applications of scanning thermal microscopy (sthm), Advanced Functional Materials 30, 1900892 (2020).
- Minnich et al. [2011] A. J. Minnich, J. A. Johnson, A. J. Schmidt, K. Esfarjani, M. S. Dresselhaus, K. A. Nelson, and G. Chen, Thermal conductivity spectroscopy technique to measure phonon mean free paths, Phys. Rev. Lett. 107, 095901 (2011).
- Regner et al. [2013] K. T. Regner, D. P. Sellan, Z. Su, C. H. Amon, A. J. H. McGaughey, and J. A. Malen, Broadband phonon mean free path contributions to thermal conductivity measured using frequency domain thermoreflectance, Nature Communications 4, 1640 (2013).
- Hu et al. [2015] Y. Hu, L. Zeng, A. J. Minnich, M. S. Dresselhaus, and G. Chen, Spectral mapping of thermal conductivity through nanoscale ballistic transport, Nature Nanotechnology 10, 701 (2015).
- Tian et al. [2018] F. Tian, B. Song, X. Chen, N. K. Ravichandran, Y. Lv, K. Chen, S. Sullivan, J. Kim, Y. Zhou, T.-H. Liu, M. Goni, Z. Ding, J. Sun, G. A. G. Udalamatta Gamage, H. Sun, H. Ziyaee, S. Huyan, L. Deng, J. Zhou, A. J. Schmidt, S. Chen, C.-W. Chu, P. Y. Huang, D. Broido, L. Shi, G. Chen, and Z. Ren, Unusual high thermal conductivity in boron arsenide bulk crystals, Science 361, 582 (2018).
- Wilson et al. [2012] R. Wilson, B. A. Apgar, L. W. Martin, and D. G. Cahill, Thermoreflectance of metal transducers for optical pump-probe studies of thermal properties, Optics express 20, 28829 (2012).
- Huang et al. [2024] X. Huang, R. Anufriev, L. Jalabert, K. Watanabe, T. Taniguchi, Y. Guo, Y. Ni, S. Volz, and M. Nomura, A graphite thermal tesla valve driven by hydrodynamic phonon transport, Nature , 1 (2024).
- Brites et al. [2012] C. D. S. Brites, P. P. Lima, N. J. O. Silva, A. Millán, V. S. Amaral, F. Palacio, and L. D. Carlos, Thermometry at the nanoscale, Nanoscale 4, 4799 (2012).
- Quintanilla and Liz-Marzán [2018] M. Quintanilla and L. M. Liz-Marzán, Guiding Rules for Selecting a Nanothermometer, Nano Today 19, 126 (2018).
- Balandin et al. [2008] A. A. Balandin, S. Ghosh, W. Bao, I. Calizo, D. Teweldebrhan, F. Miao, and C. N. Lau, Superior Thermal Conductivity of Single-Layer Graphene, Nano Letters 8, 902 (2008).
- Ghosh et al. [2008] S. Ghosh, I. Calizo, D. Teweldebrhan, E. P. Pokatilov, D. L. Nika, A. A. Balandin, W. Bao, F. Miao, and C. N. Lau, Extremely high thermal conductivity of graphene: Prospects for thermal management applications in nanoelectronic circuits, Applied Physics Letters 92, 151911 (2008).
- Braun et al. [2022] O. Braun, R. Furrer, P. Butti, K. Thodkar, I. Shorubalko, I. Zardo, M. Calame, and M. L. Perrin, Spatially mapping thermal transport in graphene by an opto-thermal method, npj 2D Materials and Applications 6, 1 (2022).
- Elhajhasan et al. [2023] M. Elhajhasan, W. Seemann, K. Dudde, D. Vaske, G. Callsen, I. Rousseau, T. F. Weatherley, J.-F. Carlin, R. Butté, N. Grandjean, et al., Optical and thermal characterization of a group-iii nitride semiconductor membrane by microphotoluminescence spectroscopy and raman thermometry, Physical Review B 108, 235313 (2023).
- Doherty et al. [2013] M. W. Doherty, N. B. Manson, P. Delaney, F. Jelezko, J. Wrachtrup, and L. C. L. Hollenberg, The nitrogen-vacancy colour centre in diamond, Physics Reports 528, 1 (2013).
- Neumann et al. [2013] P. Neumann, I. Jakobi, F. Dolde, C. Burk, R. Reuter, G. Waldherr, J. Honert, T. Wolf, A. Brunner, J. H. Shim, D. Suter, H. Sumiya, J. Isoya, and J. Wrachtrup, High-precision nanoscale temperature sensing using single defects in diamond, Nano Letters 13, 2738 (2013).
- Toyli et al. [2013] D. M. Toyli, C. F. d. l. Casas, D. J. Christle, V. V. Dobrovitski, and D. D. Awschalom, Fluorescence thermometry enhanced by the quantum coherence of single spins in diamond, Proceedings of the National Academy of Sciences 110, 8417 (2013).
- Kucsko et al. [2013] G. Kucsko, P. C. Maurer, N. Y. Yao, M. Kubo, H. J. Noh, P. K. Lo, H. Park, and M. D. Lukin, Nanometre-scale thermometry in a living cell, Nature 500, 54 (2013).
- Laraoui et al. [2015] A. Laraoui, H. Aycock-Rizzo, Y. Gao, X. Lu, E. Riedo, and C. A. Meriles, Imaging thermal conductivity with nanoscale resolution using a scanning spin probe, Nature Communications 6, 8954 (2015).
- Acosta et al. [2013] V. M. Acosta, E. Bauch, M. P. Ledbetter, A. Waxman, L.-S. Bouchard, and D. Budker, Temperature dependence of the nitrogen-vacancy magnetic resonance in diamond, Physical Review Letters 104, 070801 (2013).
- Barry et al. [2020] J. F. Barry, J. M. Schloss, E. Bauch, M. J. Turner, C. A. Hart, L. M. Pham, and R. L. Walsworth, Sensitivity optimization for nv-diamond magnetometry, Rev. Mod. Phys. 92, 015004 (2020).
- Wojciechowski et al. [2018] A. M. Wojciechowski, M. Karadas, C. Osterkamp, S. Jankuhn, J. Meijer, F. Jelezko, A. Huck, and U. L. Andersen, Precision temperature sensing in the presence of magnetic field noise and vice-versa using nitrogen-vacancy centers in diamond, Applied Physics Letters 113, 013502 (2018).
- Gottscholl et al. [2021] A. Gottscholl, M. Diez, V. Soltamov, C. Kasper, D. Krauße, A. Sperlich, M. Kianinia, C. Bradac, I. Aharonovich, and V. Dyakonov, Spin defects in hBN as promising temperature, pressure and magnetic field quantum sensors, Nature Communications 12, 4480 (2021).
- Beaufils et al. [2018] C. Beaufils, W. Redjem, E. Rousseau, V. Jacques, A. Y. Kuznetsov, C. Raynaud, C. Voisin, A. Benali, T. Herzig, S. Pezzagna, J. Meijer, M. Abbarchi, and G. Cassabois, Optical properties of an ensemble of g-centers in silicon, Physical Review B 97, 035303 (2018).
- Kraus et al. [2014] H. Kraus, V. A. Soltamov, F. Fuchs, D. Simin, A. Sperlich, P. G. Baranov, G. V. Astakhov, and V. Dyakonov, Magnetic field and temperature sensing with atomic-scale spin defects in silicon carbide, Scientific Reports 4, 5303 (2014).
- Anisimov et al. [2016] A. N. Anisimov, D. Simin, V. A. Soltamov, S. P. Lebedev, P. G. Baranov, G. V. Astakhov, and V. Dyakonov, Optical thermometry based on level anticrossing in silicon carbide, Scientific Reports 6, 33301 (2016).
- Zhou et al. [2017] Y. Zhou, J. Wang, X. Zhang, K. Li, J. Cai, and W. Gao, Self-protected thermometry with infrared photons and defect spins in silicon carbide, Physical Review Applied 8, 044015 (2017).
- [50] See Supplemental Material at … for further details on experimental and theoretical methods and additional data.
- Zhang et al. [2021] S.-C. Zhang, Y. Dong, B. Du, H.-B. Lin, S. Li, W. Zhu, G.-Z. Wang, X.-D. Chen, G.-C. Guo, and F.-W. Sun, A robust fiber-based quantum thermometer coupled with nitrogen-vacancy centers, Review of Scientific Instruments 92, 044904 (2021).
- Cambria et al. [2023] M. C. Cambria, G. Thiering, A. Norambuena, H. T. Dinani, A. Gardill, I. Kemeny, V. Lordi, A. Gali, J. R. Maze, and S. Kolkowitz, Physically motivated analytical expression for the temperature dependence of the zero-field splitting of the nitrogen-vacancy center in diamond, Phys. Rev. B 108, L180102 (2023).
- Li et al. [2012] W. Li, N. Mingo, L. Lindsay, D. A. Broido, D. A. Stewart, and N. A. Katcho, Thermal conductivity of diamond nanowires from first principles, Physical Review B—Condensed Matter and Materials Physics 85, 195436 (2012).
- Fugallo et al. [2013] G. Fugallo, M. Lazzeri, L. Paulatto, and F. Mauri, Ab initio variational approach for evaluating lattice thermal conductivity, Phys. Rev. B 88, 045430 (2013).
- Swinkels et al. [2015] M. v. Swinkels, M. Van Delft, D. Oliveira, A. Cavalli, I. v. Zardo, R. Van Der Heijden, and E. Bakkers, Diameter dependence of the thermal conductivity of inas nanowires, Nanotechnology 26, 385401 (2015).
- Martin et al. [2009] P. Martin, Z. Aksamija, E. Pop, and U. Ravaioli, Impact of phonon-surface roughness scattering on thermal conductivity of thin si nanowires, Physical review letters 102, 125503 (2009).
- Hochbaum et al. [2008] A. I. Hochbaum, R. Chen, R. D. Delgado, W. Liang, E. C. Garnett, M. Najarian, A. Majumdar, and P. Yang, Enhanced thermoelectric performance of rough silicon nanowires, Nature 451, 163 (2008).
- Li et al. [2003] D. Li, Y. Wu, P. Kim, L. Shi, P. Yang, and A. Majumdar, Thermal conductivity of individual silicon nanowires, Applied Physics Letters 83, 2934 (2003).
- Maire et al. [2017] J. Maire, R. Anufriev, and M. Nomura, Ballistic thermal transport in silicon nanowires, Scientific Reports 7, 41794 (2017).
- Ziman [2001] J. M. Ziman, Electrons and phonons: the theory of transport phenomena in solids (Oxford university press, 2001).
- Simoncelli et al. [2020] M. Simoncelli, N. Marzari, and A. Cepellotti, Generalization of fourier’s law into viscous heat equations, Phys. Rev. X 10, 011019 (2020).
- Dragašević and Simoncelli [2024a] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors (2024a), arXiv:2303.12777 [cond-mat].
- Gandolfi et al. [2020] M. Gandolfi, C. Giannetti, and F. Banfi, Temperonic crystal: A superlattice for temperature waves in graphene, Phys. Rev. Lett. 125, 265901 (2020).
- Raya-Moreno et al. [2022a] M. Raya-Moreno, X. Cartoixà, and J. Carrete, BTE-Barna: An extension of almaBTE for thermal simulation of devices based on 2D materials, Computer Physics Communications 281, 108504 (2022a).
- Raya-Moreno et al. [2022b] M. Raya-Moreno, J. Carrete, and X. Cartoixà, Hydrodynamic signatures in thermal transport in devices based on two-dimensional materials: An ab initio study, Physical Review B 106, 014308 (2022b).
- Li et al. [2014] W. Li, J. Carrete, N. A. Katcho, and N. Mingo, Shengbte: A solver of the boltzmann transport equation for phonons, Computer Physics Communications 185, 1747 (2014).
- Ratsifaritana and Klemens [1987] C. Ratsifaritana and P. Klemens, Scattering of phonons by vacancies, International journal of thermophysics 8, 737 (1987).
- Carruthers [1961] P. Carruthers, Theory of thermal conductivity of solids at low temperatures, Reviews of Modern Physics 33, 92 (1961).
- Note [1] We note that this is formally analogous to Bosanquet-type regularizations for the transport coefficients of rarefied fluids [75, 62].
- Cepellotti and Marzari [2017] A. Cepellotti and N. Marzari, Boltzmann transport in nanostructures as a friction effect, Nano Letters 17, 4675 (2017).
- Maassen and Lundstrom [2015] J. Maassen and M. Lundstrom, Steady-state heat transport: Ballistic-to-diffusive with fourier’s law, Journal of Applied Physics 117 (2015).
- Anthony et al. [1991] T. R. Anthony, J. L. Fleischer, J. R. Olson, and D. G. Cahill, The thermal conductivity of isotopically enriched polycrystalline diamond films, Journal of Applied Physics 69, 8122 (1991).
- Tian et al. [2024] S. Tian, T. Wang, H. Chen, D. Ma, and L. Zhang, Phonon coherent transport leads to an anomalous boundary effect on the thermal conductivity of a rough graphene nanoribbon, Phys. Rev. Appl. 21, 064005 (2024).
- Polanco et al. [2023] C. A. Polanco, A. van Roekeghem, B. Brisuda, L. Saminadayar, O. Bourgeois, and N. Mingo, The phonon quantum of thermal conductance: Are simulations and measurements estimating the same quantity?, Science Advances 9, eadi7439 (2023).
- Michalis et al. [2010] V. K. Michalis, A. N. Kalarakis, E. D. Skouras, and V. N. Burganos, Rarefaction effects on gas viscosity in the Knudsen transition regime, Microfluidics and Nanofluidics 9, 847 (2010).
- Chambers [1950] R. Chambers, The conductivity of thin wires in a magnetic field, Proceedings of the Royal Society of London. Series A. Mathematical and Physical Sciences 202, 378 (1950).
- Mi et al. [2019] S. Mi, A. Toros, T. Graziosi, and N. Quack, Non-contact polishing of single crystal diamond by ion beam etching, Diamond and Related Materials 92, 248 (2019).
- Burek et al. [2014] M. J. Burek, Y. Chu, M. S. Z. Liddy, P. Patel, J. Rochman, S. Meesala, W. Hong, Q. Quan, M. D. Lukin, and M. Lončar, High quality-factor optical nanocavities in bulk single-crystal diamond, Nature Communications 5, 5718 (2014).
- Graziosi et al. [2018] T. Graziosi, S. Mi, M. Kiss, and N. Quack, Single crystal diamond micro-disk resonators by focused ion beam milling, Apl Photonics 3 (2018).
- Riedrich-Möller et al. [2012] J. Riedrich-Möller, L. Kipfstuhl, C. Hepp, E. Neu, C. Pauly, F. Mücklich, A. Baur, M. Wandt, S. Wolff, M. Fischer, S. Gsell, M. Schreck, and C. Becher, One- and two-dimensional photonic crystal microcavities in single crystal diamond, Nature Nanotechnology 7, 69 (2012).
- Cai et al. [2019] Y.-Y. Cai, E. Sung, R. Zhang, L. J. Tauzin, J. G. Liu, B. Ostovar, Y. Zhang, W.-S. Chang, P. Nordlander, and S. Link, Anti-stokes emission from hot carriers in gold nanorods, Nano Letters 19, 1067 (2019), publisher: American Chemical Society.
- Carattino et al. [2018] A. Carattino, M. Caldarola, and M. Orrit, Gold Nanoparticles as Absolute Nanothermometers, Nano Letters 18, 874 (2018).
- Jollans et al. [2020] T. Jollans, M. Caldarola, Y. Sivan, and M. Orrit, Effective Electron Temperature Measurement Using Time-Resolved Anti-Stokes Photoluminescence, The Journal of Physical Chemistry A 124, 6968 (2020).
- Crampton et al. [2018] K. T. Crampton, A. Fast, E. O. Potma, and V. A. Apkarian, Junction Plasmon Driven Population Inversion of Molecular Vibrations: A Picosecond Surface-Enhanced Raman Spectroscopy Study, Nano Letters 18, 5791 (2018).
- Dubi and Sivan [2019] Y. Dubi and Y. Sivan, “hot” electrons in metallic nanostructures—non-thermal carriers or heating?, Light: Science & Applications 8, 89 (2019).
- Ziegler [1985] J. P. Ziegler, James F.and Biersack, The stopping and range of ions in matter, in Treatise on Heavy-Ion Science, Vol. 6, edited by D. A. Bromley (Springer US, Boston, MA, 1985) pp. 93–129.
- Peierls [1929] R. Peierls, Zur kinetischen theorie der wärmeleitung in kristallen, Annalen der Physik 395, 1055 (1929).
- Peierls [1996] R. E. Peierls, Quantum theory of solids (Clarendon Press, 1996).
- Simoncelli et al. [2019] M. Simoncelli, N. Marzari, and F. Mauri, Unified theory of thermal transport in crystals and glasses, Nature Physics 15, 809 (2019).
- Simoncelli et al. [2022] M. Simoncelli, N. Marzari, and F. Mauri, Wigner formulation of thermal transport in solids, Physical Review X 12, 041011 (2022).
- Di Lucente et al. [2023] E. Di Lucente, M. Simoncelli, and N. Marzari, Crossover from boltzmann to wigner thermal transport in thermoelectric skutterudites, Physical Review Research 5, 033125 (2023).
- Cepellotti and Marzari [2016] A. Cepellotti and N. Marzari, Thermal transport in crystals as a kinetic theory of relaxons, Physical Review X 6, 041013 (2016).
- Dragašević and Simoncelli [2024b] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors (2024b), arXiv:2303.12777 [cond-mat.mtrl-sci] .
- Note [2] In particular, Eq. 6 of Ref. [53] encompasses as special cases both thin films and thin wires, as discussed after Eq. 5 of Ref. [76] and Eq. 11.2.9 of Ref. [60].
- Dragašević and Simoncelli [2023] J. Dragašević and M. Simoncelli, Viscous heat backflow and temperature resonances in extreme thermal conductors, arXiv preprint arXiv:2303.12777 (2023).
- Mathematica [2019] Mathematica, Wolfram Research Inc., Champaign, IL (Version 12.0, 2019).
- Allen and Perebeinos [2018] P. B. Allen and V. Perebeinos, Temperature in a peierls-boltzmann treatment of nonlocal phonon heat transport, Physical Review B 98, 085427 (2018).
- Sendra et al. [2022] L. Sendra, A. Beardo, J. Bafaluy, P. Torres, F. X. Alvarez, and J. Camacho, Hydrodynamic heat transport in dielectric crystals in the collective limit and the drifting/driftless velocity conundrum, Physical Review B 106, 155301 (2022).
- Gürcan et al. [2013] Ö. D. Gürcan, P. Diamond, X. Garbet, V. Berionni, G. Dif-Pradalier, P. Hennequin, P. Morel, Y. Kosuga, and L. Vermare, Transport of radial heat flux and second sound in fusion plasmas, Physics of Plasmas 20 (2013).
I Supplementary material
I.1 Sample preparation
We use commercial (100) CVD single crystalline diamond substrate (Hiqute Diamond, Rouge-T12-100), enriched to above 99.95% in 12C isotopes and containing an NV-rich layer (1–2 ppm) of µm thickness at the surface. We first perform preliminary acid cleaning: piranha solution (H2SO4:H2O2 3:1) and hydrofluoric acid remove organics and metallic residuals, respectively. The substrate is then mounted on a silicon chip with 2 µm thick wet oxide via QuickStick wax. We apply a reported non-contact ion beam polishing method to improve the surface roughness [77].
To define the chromium patches, a 150 nm-thick Cr layer is evaporated on the NV-rich diamond surface. The patches are then defined using electron beam lithography followed by wet etching. 150 nm of hydrogen silsesquioxane (HSQ; XR-1541-006) is spin-coated, written with electron beam lithography, and subsequently developed in tetramethylammonium hydroxide (TMAH 25) as a mask. TechniEtch Cr01 from MicroChemicals ((NH4)2Ce(NO3)6+HClO4) removes the unmasked Cr layer. Non-contact ion polishing is used to clean the sample surface, and then the patches are released by HF 1%.
We then define the cantilevers, aligned to the Cr patches, using a second round of electron beam lithography followed by several etching steps. A thin titanium layer (40 nm) is evaporated on the surface. A 2 µm-thick layer of HSQ (FOx-16) is spin coated above and exposed with electron beam lithography followed by the aforementioned development process. The Ti layer serves as a conductive and adhesive layer between diamond and HSQ. A Chlorine-based reactive ion etching process is used to pattern the Ti layer (STS Multiplex ICP, 800 W ICP power, 150 W bias power, 10 sccm Cl2 and 10 sccm BCl3, 3 mTorr). To transfer the pattern from the HSQ hard mask to the diamond, a vertical etch is applied with directional oxygen plasma (STS Multiplex ICP, 400 W ICP power, 200 W bias power, 50 sccm O2, 15 mTorr). Faraday cage angled etch is then used under the same plasma condition as the previous step to undercut the structures [78]. The hard mask is stripped in the HF 1% bath.
Since the Faraday cage angled etch technique restricts the lateral dimensions of suspended structures, we proceed further with focused ion beam (FIB) milling. This method allows greater flexibility in structuring, enabling us to precisely polish narrow cantilevers and undercut wider ones [79, 80]. A 200 nm Ti layer is deposited to shield the sample from gallium (Ga) implantation and to prevent drift during milling caused by charge accumulation. The milling is conducted with a Zeiss CrossBeam 540 SEM/FIB dual beams system using 30 kV Ga+ ions on both sides of each cantilever. We apply cross-section mode with 1.5 nA beam current to sputter away most of the material and lower the ion current to 700 pA to realize better sidewall smoothness. Finally, cantilevers with finely polished sidewalls are revealed after removing the Ti layer in HF bath, as shown in Fig. 1b.
I.2 Experimental setup
All experiments are carried out in our home-built confocal microscope. The diamond sample is mounted on a temperature-controlled stage, kept at a constant temperature of 22 ∘C (except for calibration measurements in Fig. 1). The output beam of a 515 nm laser is split into two beams, one for heating the chromium patch at the apex of the cantilever and the other one for reading out the NV centers’ fluorescence all along the cantilever. The power of each beam can be controlled independently. The two beams are focused on different positions of the cantilevers with a high numerical aperture microscope objective (NA=0.9). The fluorescence is collected into a multi-mode fiber and detected with a multi-pixel photon counter (Hamamatsu C14455-1550GA).
Microwaves are delivered to the sample with a wire loop placed in close vicinity to the diamond sample. The microwave frequency is sinusoidally modulated at 5 kHz by the microwave generator (Rohde&Schwarz SMB 100A) to be used in combination with the lock-in amplifier. The readout parameters are chosen to optimize temperature sensitivity: the readout laser power is 0.15 mW, the microwave power is 41 dBm and the microwave frequency modulation depth is 5 MHz (corresponding to approximately half of the ODMR linewidth).
Fig. S2 shows how the above parameters affect the measured value of the zero-field splitting (ZFS). We find that, on cantilevers fabricated with FIB milling, the readout laser induces a small heating (blue squares). On the other hand, no laser-induced heating is measured on cantilevers fabricated only with Faraday cage angle etch, without the final FIB milling step (red triangles). We conclude that the layer contaminated with Ga ions during milling causes absorption of the readout laser light, which in turn generates local heating. Nevertheless, the small, K, heating caused by the readout laser with 0.15 mW power is not expected to affect the temperature profile measured in the presence of the heating laser on the metal patch: this contribution from the readout laser is removed when subtracting the ZFS values without the heating laser (see Fig. S8). The microwave power is found to have a negligible effect on the temperature readout. Note that no static magnetic field is applied to the sample. This magnetic field-free scheme to measure the ZFS for thermometry is known to be insensitive to small magnetic field fluctuations [51].
I.3 Independent study of the patch temperature
To confirm that the metal patch is well thermalized with the diamond cantilever, we measure the hot carrier anti-Stokes emission of the patch to get an independent estimate of its temperature. Luminescence in noble metals following photoexcitation is a complex process influenced by various mechanisms, including electron-electron scattering, electron-phonon coupling, multiphoton excitations, and surface plasmons [81, 82]. Anti-Stokes emission corresponds to photons emitted at a higher energy than the energy of the laser incident on the material. As detailed in previous studies, the anti-Stokes emission spectrum can be fit with Boltzmann statistics to obtain a first approximation of the effective temperature [81, 83]. Under the assumption that both the photonic and electronic densities of states vary much slower than the Boltzmann factor close to the laser energy [84, 83], we have the simple expression for the anti-Stokes intensity as a function of the Raman shift
from which we can determine the effective temperature .
We measure the anti-Stokes emission spectrum on three different cantilevers, using 532 nm continuous wave excitation and a spectrometer. A notch filter rejects most of the reflected excitation laser. Collected spectra are shown in Fig. S3a and can be well fitted with the exponential decay from Eq. (S1) over the range that we can experimentally access. Note that we discard the energy window corresponding to the Raman peak of diamond, to avoid spurious residual contribution from the diamond cantilever. We extract the anti-Stokes temperature from the fit. This experiment was performed on a patch made of gold (Au) instead of Cr, but with a few nm-thin Cr layer used as adhesion layer between Au and diamond. Hence, aside from the absorption coefficient of the patch , no major qualitative difference is expected in the metal-to-diamond thermalization efficiency.
The relationship between and the lattice temperature is in general not straightforward, specifically under continuous wave illumination [85]. Hence, we perform a calibration measurement to relate the extracted to the patch temperature, using the temperature-controlled stage on which the diamond sample is mounted (see inset of Fig. S3b). We find that the relation between and the controlled patch temperature can be well described using a linear fit: . We can therefore determine the temperature of metal patches versus heating laser power, on cantilevers with different widths, as shown in Fig. S3b. Although this method does not grant high precision (note also that the extracted zero-laser-power temperature is significantly lower on the 1.18 µm cantilever), this measurement proves that the metal patch temperature increases linearly when heating laser is applied. Temperature estimates with anti-Stokes emission are also found to be of the same order as values presented in the main text on diamond cantilevers for similar heating power, meaning that the interface thermal resistance between the patch and diamond is sufficiently small.
We can estimate convective and radiative thermal losses from the metal patch using the above temperature measurements. For the range of applied laser power, we get patch temperature . We use this value to get an upper-bound estimate of radiative and convective losses. For radiative heat transfer, we use the Stefan-Boltzmann law for black-body radiation:
where is the Stefan-Boltzmann constant and is the emissivity of the body. Even for a perfectly black body, i.e., , we get an upper bound W. Convective heat transfer can be estimated as
with the patch area and the ambient air temperature. Considering a value of the heat transfer coefficient for free convection with air (which is on the higher end of values reported in the literature), µm2 and , we get an estimated W. Both estimates for and are 5 to 6 orders of magnitude lower than the applied heating laser power. Our main conclusion is thus that convective and radiative thermal losses from the metal patch to the surrounding can safely be neglected. The same is also true for radiative and convective losses in the diamond cantilever itself (it has even lower temperature and its outer surface is only 10 to 100x bigger than the patch).
I.4 Experimental estimate of the effective thermal conductivity
The effective thermal conductivity of our cantilevers is defined as . The temperature gradient is directly extracted from the measured temperature profile. We only need to estimate the heat flux density in order to obtain . As introduced in the main text, we have . The heating laser power incident on the Cr patch is measured directly. The Cr patch absorption coefficient is estimated by measuring its reflection coefficient and assuming negligible transmission and scattering, i.e. . Finally, the cantilever cross-section is , where both the width and cross-section angle are estimated from SEM images of the different cantilevers, see for example Fig. 1b and Fig. 3a of the main text. Note that the explicit expression for as a function of the above experimental parameters, specifically and , is:
We conservatively assume an uncertainty of 50 nm in the extracted width from SEM images. This uncertainty needs to be properly taken into account in the uncertainty of , together with the uncertainty on the temperature gradient. We find that the width uncertainty is the dominant factor, explaining the significantly bigger error bars in , see Fig. 4, compared to the temperature gradient in Fig. 3c.
I.5 Additional figures
I.6 Theoretical analysis
An established approach to describe thermal transport in crystals from first principles is the linearized phonon Boltzmann transport equation (LBTE) [87, 88, 60]:
where is phonon distribution (, with being the phonon wavevector and the phonon mode index). The LBTE is a microscopic model, which describes the propagation and scattering of phonons in the presence of a temperature gradient. The LBTE can be parametrized from first principles, and its linear-response solution allows to predict the thermal conductivity. We note, in passing, that the LBTE has been recently generalized to account not only for the ability of phonons to propagate particle-like, but also for their capability to tunnel wave-like [89, 90, 91]. In diamond, propagation dominates over tunneling and the LBTE accurately describes thermal transport [61]. Moreover, since in diamond momentum-conserving “normal” phonon scattering processes have higher (or comparable) strength than momentum-non-conserving ”Umklapp” processes, a solution of the LBTE accounting for the full scattering matrix is necessary to properly capture heat-transport properties [53, 54].
The first-principles solution of the LBTE with full scattering matrix has the drawback of being computationally expensive, making it impractical to explore how nanostructure shape affects heat transport. Therefore, to investigate thermal transport in cantilevers having different cross-section shapes and sizes, we rely on the mesoscopic viscous heat equations (VHE) [61]. The VHE are two coupled partial differential equations for the mesoscopic temperature () and phonon drift-velocity () fields. They are obtained from coarse-graining the LBTE[92, 61]. The VTE encompasses both the diffusive (Fourier) and hydrodynamic regimes [7, 92] of thermal transport, as well as the crossover between them. It has been shown that the VHE rationalizes the experimental demonstration of hydrodynamic transport in graphite [61], and also yields predictions compatible with the space-dependent solution of the LBTE[64, 65] in complex-shaped devices [93]. In the steady-state regime, for a material with isotropic transport properties such as diamond, the VHE read:
where is the phonon viscosity, the coupling coefficients and arise from the energy-crystal momentum relation for phonons [61], while accounts for the presence of momentum-dissipating scattering processes [61]. We use capital to emphasize the space variation at the mesoscopic scale.
From Eq. (S6) one can derive energy and momentum balance equations in terms of a heat flux generalized to the hydrodynamic regime:
where labels Cartesian directions.
The mesoscopic heat flux is then used to evaluate the effective thermal conductivity across the cantilevers. For a cantilever, taken along axis, of length and with an imposed temperature difference at its boundaries , we get:
where is the average heat flux discussed in the main text.
All the parameters entering the VHE can be determined from the first-principles solution of the full LBTE (i.e. accounting for the actual phonon band structure and full collision matrix). An exact solution of the LBTE has been discussed for the case of nanowires [53, 66] under the assumption of diffusive boundaries. This approach allows to account for the influence of boundary scattering on the conductivity of wires, providing an expression which depends only on the cross-section of the wire 222In particular, Eq. 6 of Ref. [53] encompasses as special cases both thin films and thin wires, as discussed after Eq. 5 of Ref. [76] and Eq. 11.2.9 of Ref. [60] and can be applied to the triangular cross-section nanowire studied in this work. Importantly, the description of finite-size effects is based on the characteristic length scale of the problem, following Casimir’s treatment of boundary scattering in small devices at low temperatures [68]. In this regime, where intrinsic anharmonic phonon scattering occurs at a rate comparable to or higher than boundary scattering, the appropriate length scale to use is one-half of the smallest system dimension [70]. After verifying with the established Casimir model [54] that boundary scattering has strong effects on the conductivity but has negligible effect on the other parameters entering the VHE (including the thermal viscosity ), we accounted for nanowire boundary scattering on the conductivity only and used bulk values (calculated in Ref. [61]) for the other parameters entering in the VHE. This approximation is further justified a posteriori, since Fig. S10 shows that viscous effects are unimportant in the cases considered here.
In summary, we determine the linear-response solution of the LBTE (Eq. (S5)) for diamond nanowires using the ShengBTE package [66] and use the bulk thermal viscosity computed following the theoretical prescriptions of Ref. [95]. Then, we solve the VHE numerically (Eq. (S6)) using a finite-element solver implemented in Mathematica [96], imposing a temperature gradient K around room temperature on µm long cantilevers. Finally, we obtain the effective thermal conductivity from Eqs (S7) and (S8). There is no fitting parameter in our model.
I.7 Estimation of the magnitude of viscous hydrodynamic effects
Violations from standard diffusive thermal transport can emerge in multiple ways: when microscopic carriers achieve a mean free path comparable to the device size [97], or when momentum-conserving phonon-phonon collisions dominate over momentum-relaxing Umklapp collisions [61]. Both ballistic and hydrodynamic effects share the common underpinnings of being described by non-local relations between temperature gradient and heat flux [98, 99, 95]. The problem of distinguishing these effects is challenging and not addressed even in pioneering works, which employ phenomenologically parametrized mesoscopic models [14, 15] or simplified microscopic models that do not capture hydrodynamic physics [30]. Depending on the relative magnitude of hydrodynamic and size effects, we can distinguish different regimes [71]: (i) “ideal diffusion”, where the thermal conductivity is an intrinsic constant; (ii) “apparent diffusion”, the regime in which anomalous, non-Fourier size-heat flux relations are driven by a conductivity that depends on size due to ballistic effects; (iii) “hydrodynamics”, which originates from the presence of viscous effects dominating over momentum-dissipation effects.
To distinguish between these different regimes, we estimate the magnitude of hydrodynamic effects in the cantilevers. In the limit of negligible hydrodynamic effects, Eqs. (S6) reduce to :
In the absence of ballistic effects, Eq. S9 is exactly Fourier’s law of diffusive heat conduction. In the presence of ballistic effects, the thermal conductivity can no longer be considered an intrinsic material property and depends on size, corresponding to the apparent diffusive regime. In this regime, the heat flux becomes
The importance of viscous effects can thus be estimated by comparing the results of computing the heat flux with Eq. (S7) (full solution) versus Eq. (S10) (apparent diffusive case). Fig. S10 shows the effective thermal conductivity computed in both cases, together with experimental data presented in the main text. The two curves show a minor difference over the whole range of width. This means that viscous effects only play a minor role in the observed deviations from pure diffusion. Ballistic effects are thus the main factor in the reduction of for reducing in 1-dimensional diamond cantilevers.
I.8 Influence of vacancy impurities